### v1

# Uplink-Downlink Duality for Beamforming in Integrated Sensing and Communications

Abstract
This paper considers the beamforming and power optimization problem for a class of integrated sensing and communications (ISAC) problems that utilize the communication signals simultaneously for sensing. We formulate the problem of minimizing the Bayesian Cramér-Rao bound (BCRB) on the mean-squared error of estimating a vector of parameters, while satisfying downlink signal-to-interference-and-noise-ratio constraints for a set of communication users at the same time. The proposed optimization framework comprises two key new ingredients. First, we show that the BCRB minimization problem corresponds to maximizing beamforming power along certain sensing directions of interest. Second, the classical uplink-downlink duality for multiple-input multiple-output communications can be extended to the ISAC setting, but unlike the classical communication problem, the dual uplink problem for ISAC may entail negative noise power and needs to include an extra condition on the uplink beamformers. This new duality theory opens doors for efficient iterative algorithm for optimizing power and beamformers for ISAC.
Index Terms:
Bayesian Cramér-Rao bound, beamforming, integrated sensing and communications, uplink-downlink duality
I Introduction
Traditional wireless communication systems focus on the task of transmitting information. Although these traditional systems typically also perform sensing-like tasks (e.g., channel estimation), the primary purpose of these sensing operations is to enable or to enhance communications. In recent years, there has been a growing interest in designing wireless systems that are capable of delivering sensing functionalities separately from and in addition to communications services [2, 3]. Such systems are expected to play a key role in emerging applications such as autonomous driving, smart cities, and industrial automation.
Moreover, there is also an emerging trend to integrate sensing operations into the communication system design. As opposed to implementing separate radar-like waveforms to perform sensing tasks independently of the communication tasks, there is now an increasing interest in exploring whether it is possible to use the same waveform to perform both sensing and communication tasks simultaneously. This brings operational benefits such as shared utilization of the wireless spectrum and hardware resources, as well as lower cost.
This paper focuses on the task of designing common waveforms for integrated sensing and communications (ISAC). This is a more challenging task than designing waveforms for communications alone, because the same information-bearing signals must also perform sensing. In the context of multiple-input multiple-output (MIMO) transmit beamforming, the beamformers that need to guarantee certain quality-of-service metrics for communication users must now also be optimized for the relevant performance metric for sensing [4, 5, 6].
Specifically, this paper considers the problem of designing an optimal set of downlink beamformers that can be used by the base station to deliver optimal sensing performance for estimating a vector of unknown parameters, while simultaneously satisfying a set of signal-to-interference-and-noise-ratio (SINR) constraints for the communications users.
To quantify the sensing performance, we utilize the Cramér-Rao bound (CRB) as a metric for sensing. The optimization of CRB is complicated due to the fact that the structure of the CRB involves a matrix inverse (e.g., see [6, 7, 8] for the scalar case where the CRB is a fraction). One of the main contributions of this paper is that, under certain conditions, the minimization of CRB can be transformed into a sequence of subproblems, each involving maximizing power in certain sensing directions of interest. This essentially reduces the ISAC problem to downlink beamforming problems that resemble those treated in classical works in MIMO communications [9, 10, 11, 12, 13, 14]. These subproblems are indexed by a set of auxiliary variables, which can be optimized in an outer loop.
Furthermore, this paper shows that an uplink-downlink duality relationship can be developed for each of these subproblems, so that the optimal downlink beamformers can be obtained by solving a dual uplink problem. But unlike the duality relationship in the classical MIMO communication setting [9, 10, 11, 12, 13, 14], the dual uplink problem for ISAC may entail negative noise power and involve extra conditions on the uplink beamformers. This paper provides a theoretical development of this new duality relationship, which can open doors for efficient numerical algorithms for power and beamforming optimization for ISAC systems.
I-A Related Works
Much existing work on ISAC has been devoted to the problem of beamforming design for simultaneous communications and sensing. In the existing literature, the beamforming design can take on one of many possible formulations: i) optimizing the sensing performance subject to a set of constraints on the user SINRs [4, 5, 6, 15, 16, 17, 18]; ii) optimizing the user rates (e.g., sum or min rate) subject to a constraint on the sensing performance metric [19, 18, 20]; or iii) optimizing a weighted average of the communication and sensing metrics [21]. The question of which formulation is more appropriate depends on the specific application. In addition, the above works also differ based on the nature of sensing operation (i.e., detection versus estimation) and the adopted performance metric. Common metrics include beampattern-based designs [4, 5, 19], classical CRB [6, 15], Bayesian CRB (BCRB) [7], radar signal-to-noise-ratio [16, 20], radar SINR [17], and radar signal-to-clutter-and-noise-ratio [18], etc.
In terms of beamformer design for ISAC, the works [4, 5, 6, 19, 7, 16, 18] all adopt the semidefinite relaxation (SDR) method, which relaxes the beamformer design problem into optimization of transmit covariance matrices. In addition to the high computational burden, these works assume that the set of beamformers can be extended to include additional beamformers dedicated to sensing only (see [4] for details). Thus, these works do not exclusively rely on the communications beamformers for sensing, in contrast to the setting of the current paper.
In a separate line of work, [17, 20, 15] develop low-complexity iterative algorithms based on successive convex approximation (SCA), which allows the beamforming problem to be solved as a sequence of second-order cone programs (SOCPs). While these algorithms tend to be more efficient than the SDR methods, it is unknown whether they guarantee global convergence. It is worthwhile to remark that there are also other suboptimal low-complexity algorithms developed for different contexts, e.g., robust beamforming for ISAC [22], and multi-cell ISAC [23].
The above works primarily design the beamformer based on average statistics for communications. A recent line of work [24, 25] has considered data-dependent beamforming in order to mitigate the effect of randomness in the transmitted signal on the sensing performance. Although [24, 25] demonstrate that such a strategy may improve the sensing, it entails a more complicated optimization; also, the performance improvement diminishes at large block lengths.
This paper advocates the use of uplink-downlink duality for beamformer design. Uplink-downlink duality is a beamforming design concept rooted in classical MIMO communications. The main idea first appears in [10], which relies on transforming the difficult downlink communication problem into an equivalent virtual uplink problem [9]. The uplink-downlink duality concept has been instrumental in the development of capacity results for the MIMO downlink channel [26, 27, 28]. Moreover, the work [12] shows that this duality concept is intimately related to the Lagrange duality theory in optimization, and it can be further extended to solving the downlink beamforming problem with per-antenna power constraints. Since then, the duality principle has been adopted to tackle the downlink beamforming design in many contexts, including multi-cell coordinated beamforming [29, 30], cooperative cellular systems [31, 32], and energy harvesting [33].
I-B Main Contributions
This paper develops an uplink-downlink duality relationship for solving an ISAC problem of minimizing CRB for sensing subject to SINR constraints for the communication users. The main contributions of this paper are as follows:
  1. 1.
    To address the issue that the CRB has a complicated fractional structure, this paper introduces a set of auxiliary variables and shows that under certain conditions, the downlink ISAC problem can be transformed into a max-min formulation with a particular quadratic objective. Such a transformation not only yields an optimization problem in a considerably more tractable form, but also gives rise to a novel interpretation that optimizing the CRB is equivalent to the maximization of beamforming power along certain spatial directions.
  2. 2.
    The classical duality relationship between the uplink and the downlink can be used to solve the downlink beamforming problem for minimizing the base station total transmit power subject to SINR constraints for the communication users. This paper shows that the ISAC beamforming problem can, in contrast, be regarded as a maximization of directional power, under a total power constraint and the SINR constraints. We develop a duality theory for this new setting, which allows the ISAC problem to be viewed as a virtual uplink communication problem with a modified noise covariance that encodes the sensing metric. Further, the uplink noise covariance is not necessarily positive semidefinite (PSD). We show that duality continues to hold under an admissibility condition. Moreover, the dual uplink problem entails an additional M-matrix constraint on the beamformers. This new duality theory opens doors for efficient numerical algorithms for beamformer design in ISAC.
I-C Paper Organization and Notation
The rest of the paper is organized as follows. Section II discusses the ISAC system model and presents the problem formulation. In Section III, we transform the ISAC problem into an equivalent max-min problem and provide interpretations of the new formulation. Section IV presents the uplink-downlink duality theory for the ISAC system. Section V presents numerical results. The paper concludes in Section VI.
This paper uses lowercase, lowercase boldface, and uppercase boldface letters to denote scalars, vectors, and matrices, respectively. We use $(\cdot)^{\textsf{T}}$, $(\cdot)^{\textsf{H}}$, $(\cdot)^{-1}$, and $(\cdot)^{\dagger}$ to represent the transpose, Hermitian, inverse, and pseudo-inverse of a matrix. We use $\Re\{\cdot\}$, $\Im\{\cdot\}$ to denote the real and imaginary parts of a complex number, $\overline{(\cdot)}$ as the complex conjugate, and $\imath$ as $\sqrt{-1}$. For vectors, $\|\cdot\|$ denotes the Euclidean norm. The matrix operators $\operatorname*{Tr}(\cdot)$ and $\rho_{\text{max}}(\cdot)$ denote the trace and spectral radius. Note that for a PSD matrix, the spectral radius is its maximum eigenvalue. We use $\mathbf{I}$ to denote an identity matrix of appropriate dimensions with its $\ell$-th column denoted by $\mathbf{e}_{\ell}$. We use $\mathbf{1}$ to denote the vector of all ones and $\mathbf{0}$ a matrix (or vector) of all zeros with appropriate dimension. We use $\operatorname*{diag}(d_{1},\ldots,d_{K})$ to denote a diagonal matrix with diagonal elements $d_{1},\ldots,d_{K}$. The space of $n\times m$ matrices with real (complex) entries is denoted by $\mathbb{R}^{n\times m}$ ($\mathbb{C}^{n\times m}$). We use $\mathbb{E}[\cdot]$ to denote the expectation of a random variable. Finally, $\mathcal{CN}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ represents the circularly-symmetric complex Gaussian vector with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$.
II System Model
Figure 1: The ISAC system model where the BS serves $K$ communication users via linear beamforming while simultaneously aiming to learn some underlying parameter of the sensing environment using $N_{\text{T}}$ transmit (Tx) antennas and $N_{\text{R}}$ receive (Rx) antennas in full-duplex operation.
This paper considers a narrowband downlink multi-user ISAC system illustrated in Figure 1, comprising a MIMO base station (BS) with $N_{\text{T}}$ transmit antennas and $N_{\text{R}}$ receive antennas. The BS provides two distinct functionalities: (i) downlink communications to $K$ single-antenna users, and (ii) monostatic sensing for estimating a vector of $L$ real-valued parameters denoted by $\boldsymbol{\eta}\in\mathbb{R}^{L}$.
We study the ISAC operation over a coherence interval spanning $T$ symbol periods in which both the sensing parameters and communication channels are assumed to be fixed. In each symbol period, the BS transmits a narrowband waveform through its transmit array that conveys communication symbols to the remote users, while collecting the back-propagated echo signal through the receive array for sensing the parameters of interest. The communications and sensing take place at different time scales. The former happens during each symbol period, whereas the latter takes place over multiple symbols within the coherence interval.
The parameters that can be estimated in this setting are essentially spatial parameters of the channel (e.g., angles of arrival in azimuth and elevation for multiple targets of interest) as opposed to temporal parameters (e.g., delay and Doppler) [^1] The estimation of delay and Doppler requires either larger bandwidth (beyond the narrowband assumption) or longer observation windows (beyond the coherence interval), and is outside of the scope of the current investigation..
II-A Transmit Signal Model
This paper adopts a signaling scheme in which the transmit vector across the antennas in the $t$-th symbol period, $\mathbf{x}[t]\in\mathbb{C}^{N_{\text{T}}}$, follows a linear beamforming model as the sum of $K$ beamformed communication symbols for the $K$ users:
$\mathbf{x}[t]=\mathbf{V}\mathbf{s}[t]=\sum_{k=1}^{K}\mathbf{v}_{k}s_{k}[t],\quad t=1,\ldots,T,$ (1)
where $\mathbf{v}_{k}$ is the beamforming vector for the $k$-th user and $\mathbf{s}[t]=\left[s_{1}[t],\ldots,s_{K}[t]\right]^{\textsf{T}}\in\mathbb{C}^{K}$ is a vector of communication symbols for the $K$ users modelled as a zero-mean i.i.d. Gaussian random vector with $\mathbb{E}[\mathbf{s}[t]\mathbf{s}^{\textsf{H}}[t]]=\mathbf{I}$. We use $\mathbf{V}\triangleq\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ to denote the matrix of beamformers. A total power constraint $P$ is imposed on the transmit signal, i.e., we require $\operatorname*{Tr}\left(\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\leq P$.
The beamforming model (1) is exactly the same as that of conventional communication systems in that the total number of beamformers is equal to the number of users. The key consideration here, however, is that the communication symbols are also used for sensing, i.e., these beamformers are designed to optimize both the sensing performance and the communication performance.
The above beamforming model can be contrasted with the so-called extended model considered in many existing ISAC studies, e.g., [4, 5, 6, 8], where the $K$ beamformers are augmented with additional beamformers for sensing, i.e.,
$\mathbf{\tilde{x}}[t]=\mathbf{V}\mathbf{s}[t]+\mathbf{V}_{\text{s}}\mathbf{s}_{\text{s}}[t].$ (2)
where $\mathbf{V}_{\text{s}}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{T}}}$ is a matrix of additional beamformers (referred to as the sensing beamformers) and $\mathbf{s}_{\text{s}}[t]\in\mathbb{C}^{N_{\text{T}}}$ is typically a known pseudo-random sequence with $\mathbb{E}[\mathbf{s}_{\text{s}}[t]\mathbf{s}^{\textsf{H}}_{\text{s}}[t]]=\mathbf{I}$. This latter model can potentially enhance the sensing performance as compared to the model in (1), since it offers additional degrees of freedom, although it is also more complex to implement.
However, there are many scenarios for which the model in (1) already performs as well as the extended model. For these systems, the model (1) is much more desirable than the extended model, since the model (1) requires no modifications to the existing communication transmission scheme. This paper restricts attention to these scenarios and studies beamformer design strategies for the class of ISAC systems where the model (1) already provides the optimal sensing performance. We will define this class of ISAC systems more precisely and give sufficient conditions for these systems in a subsequent section, after first describing the respective channel models and performance metrics for sensing and communications.
II-B Sensing Model and Performance Metric
The sensing model adopted in this paper assumes that the BS can perform self-interference mitigation [34], i.e., the leakage from the transmit array to the received antennas can be cancelled. Consequently, the baseband signal signal $\mathbf{y}_{\text{s}}[t]\in\mathbb{C}^{N_{\text{R}}}$ received at the BS during the $t$-th symbol period can be written as
$\mathbf{y}_{\text{s}}[t]=\mathbf{G}^{\left(\boldsymbol{\eta}\right)}\mathbf{x}[t]+\mathbf{n}_{\text{s}}[t],\quad t=1,\ldots,T$ (3)
where $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ is the sensing channel matrix that models the “round-trip-return” channel between the BS transmit and receive antenna arrays, and $\mathbf{n}_{\text{s}}[t]\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I})$ is the additive noise at the receiver.
Observe that the sensing channel is parameterized by the parameters of interest $\boldsymbol{\eta}$. Throughout this paper, we do not specify the exact parametric dependency between $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}$ and $\boldsymbol{\eta}$. We only make the assumptions that
  • $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}$ is a deterministic function of $\boldsymbol{\eta}$, and
  • the functional form of $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}$ is known.
These are reasonable assumptions as the parametric form $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}$ can typically be obtained through appropriate channel modelling.
The goal for sensing is to estimate $\boldsymbol{\eta}$. The above model implicitly assumes that it is of interest to estimate all the parameters of the channel $\mathbf{G}^{\left(\boldsymbol{\eta}\right)}$. The scenario in which the sensing objective is to estimate only some part of $\boldsymbol{\eta}$ can be handled by incorporating a weighting factor for each component of $\boldsymbol{\eta}$ in the sensing objective, and then setting the weights corresponding to the “nuisance” parameters of no interest to zero, as will be explained in more detail shortly. In this way, the above formulation can encompass a wide variety of sensing problems, including, e.g., estimating angles-of-arrival of multiple targets or target position in the near-field.
This work adopts a Bayesian estimation framework by assuming the availability of a prior distribution $f(\boldsymbol{\eta})$ and uses the BCRB to quantify the sensing performance for estimating $\boldsymbol{\eta}$. The prior distribution $f(\boldsymbol{\eta})$ can be obtained, for example, from historical measurements in previous coherence intervals.
Given the input to the channel $\mathbf{X}\triangleq\left[\mathbf{x}[1],\ldots,\mathbf{x}[T]\right]$ and the observed output $\mathbf{Y}_{\text{s}}\triangleq[\mathbf{y}_{\text{s}}[1],\ldots,\mathbf{y}_{\text{s}}[T]]$ within the coherence interval, let $\hat{\boldsymbol{\eta}}\triangleq\hat{\boldsymbol{\eta}}\left(\mathbf{X},\mathbf{Y}_{\text{s}}\right)$ denote any well-behaved estimator of $\boldsymbol{\eta}$ [35]. The BCRB states that
$\mathbb{E}_{\boldsymbol{\eta}}\left[\mathbb{E}_{\mathbf{X},\mathbf{Y}_{\text{s}}}\left[\left(\boldsymbol{\eta}-\hat{\boldsymbol{\eta}}\right)\left(\boldsymbol{\eta}-\hat{\boldsymbol{\eta}}\right)^{\textsf{T}}\big{|}\boldsymbol{\eta}\right]\right]\succcurlyeq\mathbf{J}^{-1}_{\mathbf{V}},$ (4)
[…]
The PSD matrix $\mathbf{J}_{\mathbf{V}}$ is the $L\times L$ Bayesian Fisher information matrix (BFIM), whose $(i,j)$-th element is [35]
$\left[\mathbf{J}_{\mathbf{V}}\right]_{ij}$ $\triangleq-\mathbb{E}_{\mathbf{X},\mathbf{Y}_{\text{s}},\boldsymbol{\eta}}\left[\frac{\partial^{2}\log{f\left(\mathbf{X},\mathbf{Y}_{\text{s}},\boldsymbol{\eta}\right)}}{\partial\eta_{i}\partial\eta_{j}}\right].$ (5)
The subscript $\mathbf{V}$ here highlights the fact that the BFIM depends on the input distribution only through the beamforming matrix $\mathbf{V}$.
It can be shown [36] that for the considered model (1)-(3) with the assumption that $\mathbf{G}^{(\boldsymbol{\eta})}$ is a deterministic function of $\boldsymbol{\eta}$, the BFIM can be expressed as the sum of two components
$\mathbf{J}_{\mathbf{V}}=\mathbf{C}+\mathbf{T}_{\mathbf{V}},$ (6)
where $\mathbf{C}\in\mathbb{R}^{L\times L}$ is a PSD matrix that depends only on the prior distribution, and $\mathbf{T}_{\mathbf{V}}\in\mathbb{R}^{L\times L}$ is a PSD matrix that captures the dependency on the beamforming matrix $\mathbf{V}$.
Let $\dot{\mathbf{G}}_{i}^{(\boldsymbol{\eta})}\triangleq\tfrac{\partial\mathbf{G}^{\left(\boldsymbol{\eta}\right)}}{\partial\eta_{i}}$. The $(i,j)$-th elements of these two matrices are given by
$\mathbf{C}]_{ij}=-\mathbb{E}_{\boldsymbol{\eta}}\left[\frac{\partial^{2}\log f(\boldsymbol{\eta})}{\partial\eta_{i}\partial\eta_{j}}\right],$ (7)
and
$\mathbf{T}_{\mathbf{V}}]_{ij}=\mathbb{E}_{\boldsymbol{\eta}}\left[\frac{T}{\sigma^{2}}\operatorname*{Tr}\left(\ddot{\mathbf{G}}_{ij}^{(\boldsymbol{\eta})}\mathbf{V}\mathbf{V}^{\textsf{H}}\right)\right],$ (8)
where $\ddot{\mathbf{G}}_{ij}^{(\boldsymbol{\eta})}\triangleq{(\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i})}^{\textsf{H}}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}+{(\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j})}^{\textsf{H}}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}$.
Asymptotically, either as $\sigma^{2}\rightarrow 0$ or as $T\rightarrow\infty$, the BCRB provides a tight bound on the mean-squared error (MSE) matrix associated with the optimal (i.e., the minimum mean-squared error (MMSE)) estimator.
Given the BFIM, one can define different possible scalar functions for measuring the sensing performance, e.g., the trace-inverse of BFIM, or the logarithm of the determinant of BFIM. In this work, we consider a weighted-sum criterion
$\mu(\mathbf{V})\triangleq\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}^{-1}_{\mathbf{V}}\right)$ (9)
where $\mathbf{W}=\operatorname*{diag}\left(w_{1},\ldots,w_{L}\right)$ is a matrix whose diagonal elements indicate the relative importance for estimating the different parameters in $\boldsymbol{\eta}$. For the case where only a subset of the parameters in $\boldsymbol{\eta}$ are of interest, the weights corresponding to the nuisance parameters can be set to be zero.
We remark that the BCRB considered in this work coincides with the notion in Van Trees [35] that treats $(\mathbf{X},\mathbf{Y}_{\text{s}})$ as random observations revealed to the estimator. In the literature, tighter bounds on the MSE are also known, e.g., the Weiss-Weinstein bound [37] and the Ziv-Zakai bound [38]. Furthermore, even within the CRB family, the so-called Miller-Chang bound [39] is reported to be tighter in some scenarios. Instead of setting the weights corresponding to the nuisance parameters to be zero, the Miller-Chang technique [39] takes statistical average over the nuisance parameters in the computation of CRB. However, these potentially tighter bounds are also more difficult to compute and to optimize. In this paper, we restrict attention to the BCRB as given by the BFIM in (6)-(8).
The main advantage of using the BCRB as the sensing metric is that the BFIM has a closed-form that is relatively simple to compute as compared to the exact MSE calculation. Furthermore, it abstracts the receiver design since the bound holds for all practical estimators. Finally, we note that unlike the classical CRB used in previous works [6, 15], the BCRB does not depend on the actual value of the unknown parameter $\boldsymbol{\eta}$ and only the prior $f(\boldsymbol{\eta})$.
II-C Communication Model and Performance Metric
From the prespective of communications, upon transmitting $\mathbf{x}[t]$, each remote user receives a scalar baseband signal, denoted by $y_{k}[t]$, as given by
$y_{k}[t]=\mathbf{h}_{k}^{\textsf{H}}\mathbf{x}[t]+n_{k}[t],\quad\forall k=1,\ldots,K,$ (10)
where $\mathbf{h}_{k}\in\mathbb{C}^{N_{\text{T}}}$ is the communication channel of the $k$-th user, and $n_{k}[t]\sim\mathcal{CN}\left(0,\sigma^{2}\right)$ is the additive white Gaussian noise. In this paper, we make the assumption that the communication channel $\mathbf{H}\triangleq\left[\mathbf{h}_{1},\ldots,\mathbf{h}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ can be perfectly estimated within each coherence interval (e.g., using uplink pilots and assuming channel reciprocity), and is known perfectly at the BS.
The goal of communications is to satisfy certain achievable rate requirements for the users, which translate to constraints on the SINR at the communication receivers. Specifically, for each of the users $k=1,\ldots,K$, we impose the following SINR constraint
$\text{SINR}^{\text{DL}}_{k}(\mathbf{V})\triangleq\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},$ (11)
where $\gamma_{k}$ denote the SINR threshold for the $k$-th user. Throughout this paper, we assume that these SINR thresholds are strictly feasible.
II-D Problem Formulation
Based on the communications and sensing performance metrics discussed in the previous section, the beamforming design problem for joint sensing and communications can be formulated as that of optimizing the sensing performance subject to constraints on the communication SINRs within each coherence interval, i.e.,
$\underset{\mathbf{V}}{\mathrm{minimize}}$~~~~ $\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}^{-1}_{\mathbf{V}}\right)$ (12a)
$\mathrm{subject\ to}$~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k,$ (12b)
$\operatorname*{Tr}\left(\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\leq P.$ (12c)
Problem (12) is challenging to solve because the objective has a complicated fractional form, which is difficult to optimize as a function of the beamforming matrix $\mathbf{V}$. Furthermore, neither the objective nor the feasible set is convex in $\mathbf{V}$.
The main goal of this paper is to show that under the following key assumption, the problem (12) can be solved efficiently using an uplink-downlink duality technique.
Assumption 1
Throughout this paper, we assume that the ISAC problem with the communication channel $\mathbf{H}$, the sensing channel $\mathbf{G}^{(\boldsymbol{\eta})}$, the prior $f(\boldsymbol{\eta})$, and power constraint $P$, has the following property. Let $\mu\in\mathbb{R}$ denote the trace of weighted BCRB, and let $\boldsymbol{\gamma}\in\mathbb{R}^{K}$ denote a vector of downlink communication SINR thresholds. We assume that whenever $(\mu,\boldsymbol{\gamma})$ is achievable using the extended model (2), i.e., there exists an extended matrix $\mathbf{V}_{\text{ex}}=\left[\begin{matrix}\mathbf{V}&\mathbf{V}_{\text{s}}\end{matrix}\right]$ with a communication beamforming matrix $\mathbf{V}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ and a sensing beamformer matrix $\mathbf{V}_{\text{s}}\in\mathbf{C}^{N_{\text{T}}\times N_{\text{T}}}$ achieving
$\mu=\operatorname*{Tr}(\mathbf{W}\mathbf{J}^{-1}_{\mathbf{V}_{\text{ex}}}),$ (13)
and
$\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\mathbf{h}_{k}^{\mathsf{H}}\mathbf{V}_{\text{s}}\mathbf{V}_{\text{s}}^{\mathsf{H}}\mathbf{h}_{k}+\sigma^{2}}\geq\gamma_{k},\quad\forall k,$ (14)
the same pair $(\mu,\boldsymbol{\gamma})$ is also achievable using $K$ beamformers alone under the same power constraint $P$, i.e., using model (1) with some $\mathbf{V}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$.
In the extended model, the extra sensing beams $\mathbf{V}_{\text{s}}$ may help improve the sensing performance, but they also consume additional power and cause extra interference to the communication users. Assumption 1 states that $\mathbf{V}_{\text{s}}=\mathbf{0}$ already achieves optimal performance, in which case the model reduces to (1).
The above assumption may seem restrictive at first glance, but recent studies have established several sufficient conditions under which Assumption 1 holds. For instance, as demonstrated in [36], when the number of users $K$ is large as compared to the number of parameters to be estimated $L$, more precisely, when
$K\geq\tfrac{L(L+1)}{4},$ (15)
using $K$ beamformers is sufficient for achieving the performance of the extended beamforming model regardless of system parameters such as $N_{\text{T}}$, $N_{\text{R}}$, and channel realization. Similar conditions have also been established in several other studies [40, 41, 42] for some special cases.
It is important to note that most of these conditions are only sufficient conditions. For instance, even when $K<\tfrac{L(L+1)}{4}$, Assumption 1 may still hold, depending on the number of transmit and receive antennas, and the particular realizations of the communication channels, the sensing channel, the prior, and the total power constraint.
II-E Example
Consider an example of estimating the angle-of-departure (AoD) and the angle-of-arrival (AoA) for a target located in the far field of a full-duplex MIMO transceiver equipped with $N_{\text{T}}=N_{\text{R}}$ antennas in a monostatic setting. The sensing channel is assumed to have a single line-of-sight (LoS) path so the AoA is the same as the AoD, and the channel can be modelled as
$\mathbf{G}^{(\boldsymbol{\eta})}_{\text{AoA}}=\alpha\mathbf{A}(\theta),$ (16)
where $\alpha\in\mathbb{C}$ is the path loss coefficient and $\theta\in\mathbb{R}$ is the AoA/AoD, and $\mathbf{A}(\cdot)\in\mathbb{C}^{N_{\text{T}}\times N_{\text{R}}}$ is the combined array response given by
$\mathbf{A}(\theta)=\mathbf{a}_{\text{R}}(\theta)\mathbf{a}_{\text{T}}^{\textsf{H}}(\theta).$ (17)
For a uniform linear array of $N_{\text{T}}$ transmit antennas, we have
$\mathbf{a}_{\text{T}}(\theta)=\left[e^{\imath\pi 0\sin(\theta)},\ldots,e^{\imath\pi\left(N_{\text{T}}-1\right)\sin(\theta)}\right]^{\textsf{T}}.$
The receive array response $\mathbf{a}_{\text{R}}(\theta)$ is defined similarly.
The set of unknown parameters $\boldsymbol{\eta}=[\Re\{\alpha\},\Im\{\alpha\},\theta]^{\textsf{T}}$ describes $\mathbf{G}^{(\boldsymbol{\eta})}$ in a deterministic fashion. The goal here is to estimate $\theta$; the path loss $\alpha$ is the nuisance parameter.
Now, assume a Gaussian prior for $\alpha$ and $\theta$ with variances $\tfrac{\sigma_{\alpha}^{2}}{2T}\sigma^{2}$ and $\tfrac{\sigma_{\theta}^{2}}{2T}\sigma^{2}$, respectively, and arbitrary means. In this case, the BFIM in (6) can be shown to be
$\mathbf{J}_{\mathbf{V}}=\frac{2T}{\sigma^{2}}\left[\begin{matrix}h_{\alpha}(\mathbf{V})+\frac{1}{\sigma_{\alpha}^{2}}&0&\Re\{h_{\theta\alpha}(\mathbf{V})\}\\ 0&h_{\alpha}(\mathbf{V})+\frac{1}{\sigma_{\alpha}^{2}}&\Im\{h_{\theta\alpha}(\mathbf{V})\}\\ \Re\{h_{\theta\alpha}(\mathbf{V})\}&\Im\{h_{\theta\alpha}(\mathbf{V})\}&h_{\theta}(\mathbf{V})+\frac{1}{\sigma_{\theta}^{2}}\end{matrix}\right]$ (18)
where
$h_{\alpha}(\mathbf{V})$ $\triangleq$ $\operatorname*{Tr}\left(\mathbb{E}_{\theta}[\mathbf{A}^{\textsf{H}}(\theta)\mathbf{A}(\theta)]\mathbf{V}\mathbf{V}^{\textsf{H}}\right),$ (19)
$h_{\theta\alpha}(\mathbf{V})$ $\triangleq$ $\operatorname*{Tr}\left(\mathbb{E}_{\alpha,\theta}[\alpha\mathbf{A}^{\textsf{H}}(\theta)\dot{\mathbf{A}}(\theta)]\mathbf{V}\mathbf{V}^{\textsf{H}}\right),$ (20)
$h_{\theta}(\mathbf{V})$ $\triangleq$ $\operatorname*{Tr}\left(\mathbb{E}_{\alpha,\theta}[|\alpha|^{2}\dot{\mathbf{A}^{\textsf{H}}}(\theta)\dot{\mathbf{A}}(\theta)]\mathbf{V}\mathbf{V}^{\textsf{H}}\right),$ (21)
and $\dot{\mathbf{A}}(\theta)\triangleq\tfrac{\partial\mathbf{A}(\theta)}{\partial\theta}$. Since the goal is to estimate $\theta$ (but not $\alpha$), we set the weights to be $\mathbf{W}=\operatorname*{diag}(0,0,1)$, so that $\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}_{\mathbf{V}}^{-1}\right)=\left[\mathbf{J}_{\mathbf{V}}^{-1}\right]_{3,3}$. By utilizing the Schur complement relation
$\left[\mathbf{J}^{-1}\right]_{3,3}=(s-\mathbf{r}^{\mathsf{H}}\mathbf{S}\mathbf{r})^{-1},$ (22)
for a $3\times 3$ matrix $\mathbf{J}=\left[\begin{matrix}\mathbf{S}&\mathbf{r}\\ \mathbf{r}^{\mathsf{H}}&s\end{matrix}\right]$ with $\mathbf{S}\in\mathbb{R}^{2\times 2},\mathbf{r}\in\mathbb{R}^{2\times 1}$ and $s\in\mathbb{R}$, it can be shown [43] that problem (12) reduces to
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{maximize}}$~~ $h_{\theta}(\mathbf{V})-\frac{\left|h_{\theta\alpha}(\mathbf{V})\right|^{2}}{h_{\alpha}(\mathbf{V})+\tfrac{1}{\sigma_{\alpha}^{2}}},$ (23)
where $\mathcal{V}$ denotes the communication constraints (12b)-(12c).
As this example shows, the BCRB minimization problem is a complicated nonconvex optimization problem with a fractional structure. This paper aims to show that under Assumption 1, this problem can be solved efficiently. We now discuss the conditions under which Assumption 1 holds.
For the above example, since $L=3$, the sufficient condition (15) for Assumption 1 to hold is $K\geq 3$. But in fact, an improved version of this condition can be derived based on a result in [36], which states $K\geq\frac{d}{2}$ is already sufficient, where $d$ is the number of distinct quadratic terms in the sensing objective. In problem (23), we have $d=4$, since $h_{\theta}(\cdot)$ and $h_{\alpha}(\cdot)$ both contain one quadratic term each, and $h_{\theta\alpha}(\cdot)$ contains two, namely
$\Re\{h_{\theta\alpha}(\mathbf{V})\}$ $=\operatorname*{Tr}\left(\mathbb{E}_{\alpha,\theta}[\alpha\mathbf{A}^{\textsf{H}}(\theta)\dot{\mathbf{A}}(\theta)+\overline{\alpha}\dot{\mathbf{A}}^{\mathsf{H}}(\theta)\mathbf{A}(\theta)]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$
$\Im\{h_{\theta\alpha}(\mathbf{V})\}$ $=\operatorname*{Tr}\left(\imath\mathbb{E}_{\alpha,\theta}[\alpha\mathbf{A}^{\textsf{H}}(\theta)\dot{\mathbf{A}}(\theta)-\overline{\alpha}\dot{\mathbf{A}}^{\mathsf{H}}(\theta)\mathbf{A}(\theta)]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$
Note that quadratic terms must be of the form $\operatorname*{Tr}\left(\mathbf{Q}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)$ where $\mathbf{Q}$ is a Hermitian matrix. Thus, the real and imaginary parts of $h_{\theta\alpha}(\mathbf{V})$ need to be counted separately; see [36] for details. With $d=4$, a sufficient condition for Assumption 1 to hold for this example of AoA estimation is simply
$K\geq 2.$ (24)
Thus, for an ISAC system serving at least two users, while estimating an AoA at the same time, Assumption 1 holds, i.e., using only communication beams can already achieve the performance of the extended model with extra sensing beams.
We remark that even when $K=1$, Assumption 1 may still hold depending on the system parameters (e.g., $N_{\text{T}}$, $N_{\text{R}}$, channel realization, and the prior, etc.) For instance, if $\alpha$ and $\theta$ are independent with $\mathbb{E}\left[\alpha\right]=0$, the problem of estimating $\theta$ in (23) reduces to
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{maximize}}$~~ $h_{\theta}(\mathbf{V})$ (25)
which has a single quadratic term in the objective, as opposed to four for the general case. Using $d=1$ in the condition in [36] yields $K\geq\frac{d}{2}=\tfrac{1}{2}$. This means that Assumption 1 holds for this special case when $K=1$.
On the other hand, when the path loss has a nonzero mean and $K=1$, recent work [44] demonstrates the necessity of having one additional sensing beamformer in some cases. In this situation, removing the extra sensing beamformer would result in a performance loss, thus Assumption 1 does not hold.
II-F Existing Method and Proposed Solution Strategy
As can be seen in the above example, the BCRB optimization problem is quite complex. The goal of this paper is to identify the structure of the BCRB optimization problem (12) that would point toward an efficient solution. In prior work, an SDR strategy for the above AoA estimation example (with the caveat that the classical CRB is used in lieu of the BCRB) has already been studied [6]. This SDR technique is also applicable to the general problem (12) (as will be shown in the next section). However, solving the problem in terms of SDR can be rather inefficient. This is because it lifts the optimization space from ${N}_{\text{T}}K$ complex dimensions to $N_{\text{T}}^{2}K$ complex dimensions by defining the PSD matrices $\mathbf{R}_{k}=\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{H}}$. As a result, solving the SDR using interior-point methods would require $\mathcal{O}(N_{\text{T}}^{6}K^{3})$ operations per iteration [45]. This order of complexity is often infeasible in practical implementations, especially when $N_{\text{T}}$ is large.
In this paper, we take the SDR as a starting point and propose a further reformulation of problem (12) that lends itself towards an efficient solution. The proposed solution strategy comprises two crucial ingredients: (i) a reformulation of the BCRB minimization objective as a weighted power maximization problem in sensing directions of interest; and (ii) a connection with the following classical problem in multiuser downlink MIMO communications with no sensing
$\underset{\mathbf{V}}{\mathrm{minimize}}$~~~~~ $\sum_{k}{\mathbf{v}}_{k}^{H}\mathbf{\tilde{Q}}{\mathbf{v}}_{k}$ (26a)
$\mathrm{subject\ to}$~~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma_{\text{c}}^{2}}\geq\gamma_{k},\quad\forall k$ (26b)
where the goal is to design downlink beamformers to satisfy SINR constraints while minimizing a weighted transmit power at the BS with $\mathbf{\tilde{Q}}\succcurlyeq 0$ as the weighting matrix.
For minimizing the sum power (i.e., when $\mathbf{\tilde{Q}}=\mathbf{I}$), a duality notion has been developed in [10] that allows the downlink problem (26) to be transformed to a virtual uplink problem, a result now commonly referred to as uplink-downlink duality [10, 11, 26, 27, 12]. The key advantage of such a transformation is that the uplink problem is more computationally efficient to solve than its downlink counterpart. The duality relationship for arbitrary $\mathbf{\tilde{Q}}\succcurlyeq 0$ is developed in [12]. This paper further develops the uplink-downlink duality theory for the BCRB minimization problem. We show how the sensing objective in (12) can be transformed, and how the classical uplink-downlink duality theory can be extended for the transformed problem.
III Minimization of BCRB as Downlink Power Maximization
We begin by deriving a key result that under Assumption 1, the minimization of the BCRB in the ISAC problem can be reformulated as a power maximization problem in the sensing directions of interest.
III-A Semidefinite Relaxation
The first step is to relax the optimization problem (12) using SDR. Define $\mathbf{R}_{k}=\mathbf{v}_{k}\mathbf{v}_{k}^{\textsf{H}}$. The relaxation of the problem (12) together with the derivation of the BCRB in (6)-(8) is
$\underset{\mathbf{R}_{1},\ldots,\mathbf{R}_{K}}{\mathrm{minimize}}$~~ $\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}_{\mathbf{R}}^{-1}\right)$ (27a)
$\mathrm{subject\ to}$~ $\frac{1}{\gamma_{k}}\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}-\sum_{i\neq k}\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{i}\mathbf{h}_{k}\geq\sigma^{2}$ (27b)
$\sum_{k}\operatorname*{Tr}\left(\mathbf{R}_{k}\right)\leq P,\quad\mathbf{R}_{k}\succcurlyeq 0~~\forall k,$ (27c)
where $\mathbf{J}_{\mathbf{R}}$ denotes, with a slight abuse of notation, the BFIM matrix expressed in $\mathbf{R}_{1},\ldots,\mathbf{R}_{k}$ (i.e., with $\mathbf{V}\mathbf{V}^{\textsf{H}}$ in (8) replaced by $\sum_{k}\mathbf{R}_{k}$), and the rank-one constraints on $\mathbf{R}_{k}$ are dropped.
For general quadratic optimization problems, the SDRs do not necessarily provide a tight relaxation of the original problem, because a rank-one solution of the SDR is not guaranteed to exist. However, under Assumption 1, the SDR must be tight. We state this result as a lemma.
Lemma 1
Under Assumption 1, the ISAC problem (12) is equivalent to its SDR (27). In other words, the SDR (27) always has a rank-one solution.
Proof:
See Appendix A. ∎
The above result allows us to work with the optimization over the covariance matrices $\{\mathbf{R}_{i}\}$ instead of the beamforming vectors $\{\mathbf{V}_{i}\}$. The main advantage here is that the SDR (27) is convex. This is because $\mathbf{J}_{\mathbf{R}}$ is an affine function of $\mathbf{R}_{k}$ by (6) and the trace-inverse function is convex, so the objective function is convex in $\mathbf{R}_{k}$; also, the constraints are linear in $\mathbf{R}_{k}$. Furthermore, assuming that the SINR constraints are strictly feasible, so Slater’s condition is satisfied, the problem (27) in fact has strong duality.
Next, we reformulate the objective function in (27). A useful technique to deal with the trace of a matrix inverse is to use the following Schur complement relation [46]
$\left[\begin{matrix}\mathbf{J}_{\mathbf{R}}&\boldsymbol{e}_{\ell}\\ \boldsymbol{e}_{\ell}^{T}&d_{\ell}\end{matrix}\right]\succcurlyeq 0\ \ \Leftrightarrow\ \ d_{\ell}\geq 0,~d_{\ell}-\mathbf{e}_{\ell}^{T}\mathbf{J}_{\mathbf{R}}^{-1}\mathbf{e}_{\ell}\geq 0,$ (28)
which allows us to express the objective function (27a) as the following minimization problem
$\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}_{\mathbf{R}}^{-1}\right)=\min_{d_{1},\ldots,d_{L}}$~~ $\sum_{\ell}d_{\ell},$ (29a)
$\mathrm{s.t.~}$~~~~ $\left[\begin{matrix}\mathbf{J}_{\mathbf{R}}&\sqrt{w}_{\ell}e_{\ell}\\ \sqrt{w}_{\ell}e_{\ell}^{\mathsf{T}}&d_{\ell}\end{matrix}\right]\succcurlyeq 0,~\forall\ell.$ (29b)
This technique is first used in the earlier work on using CRB to optimize waveform for MIMO radar [47] and has since been applied extensively in radar and ISAC settings [48, 49, 6].
Now, substituting (29) into (27), we obtain
$\underset{\mathbf{R}_{1},\ldots,\mathbf{R}_{K},d_{1},\ldots,d_{L}}{\mathrm{minimize}}$~~ $\sum_{\ell}d_{\ell}$ (30a)
[…]
Problem (12) is now reformulated into a form that can be solved by a standard semidefinite programming solver. When Assumption 1 holds, this SDR must have a rank-one solution. However, as already mentioned, SDR has high complexity. In this paper, we take this SDR formulation as the starting point, and proceed by analyzing its structure by taking its Lagrangian dual. This would eventually enable algorithmic development in the beamformer domain rather than the covariance domain.
III-B Analysis of the Dual Problem
The reformulated problem (30) is convex with strong duality. We proceed by taking its Lagrangian dual with respect to (29b). Let $\tilde{\mathbf{B}}_{1},\ldots,\tilde{\mathbf{B}}_{L}\in\mathbb{R}^{(L+1)\times(L+1)}$ denote the dual variables associated with (29b)
$\tilde{\mathbf{B}}_{\ell}\triangleq\left[\begin{matrix}\mathbf{B}_{\ell}&-\boldsymbol{\beta}_{\ell}\\ -\boldsymbol{\beta}_{\ell}^{\mathsf{T}}&b_{\ell}\end{matrix}\right]\succcurlyeq 0,$ (31)
with $\mathbf{B}_{\ell}\in\mathbb{R}^{L\times L}$, $\boldsymbol{\beta}_{\ell}\in\mathbb{R}^{L}$, and $b_{\ell}\in\mathbb{R}$.
The Lagrangian dual of the problem (30) can then be written down as
$\underset{\tilde{\mathbf{B}}_{1},\ldots,\tilde{\mathbf{B}}_{L}}{\mathrm{maximize}}\min_{\underset{\{d_{\ell}\}}{\{\mathbf{R}_{k}\}\in\mathcal{R},}}\sum_{\ell}d_{\ell}(1-b_{\ell})+2\sqrt{w_{\ell}}\mathbf{e}_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\operatorname*{Tr}(\mathbf{B}_{\ell}\mathbf{J}_{\mathbf{R}})$
where $\mathcal{R}$ denotes the constraints (27b)-(27c).
Consider the inner minimization of $d_{\ell}$ for fixed $b_{\ell}$, then the outer maximization over $b_{\ell}$. It is easy to conclude that the optimal $b^{*}_{\ell}=1,\forall\ell$, as otherwise the inner problem is unbounded below. Substituting $b_{\ell}^{*}=1$, we get
$\underset{\mathbf{B}_{\ell}\succcurlyeq\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}},\forall\ell}{\mathrm{maximize}}~~\min_{\{\mathbf{R}_{k}\}\in\mathcal{R}}\sum_{\ell}2\sqrt{w_{\ell}}\mathbf{e}_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\operatorname*{Tr}(\mathbf{B}_{\ell}\mathbf{J}_{\mathbf{R}}),$ (32)
where the condition $\mathbf{B}_{\ell}\succcurlyeq\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}$ follows from $\tilde{\mathbf{B}}_{\ell}\succcurlyeq 0$ and $b_{\ell}=1$, so the Schur complement must satisfy $\mathbf{B}_{\ell}-\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\succcurlyeq 0$.
Now, since strong duality holds, the primal optimal $\mathbf{R}^{*}_{1},\ldots,\mathbf{R}^{*}_{K}$ and the dual optimal $\mathbf{B}_{1}^{*},\ldots,\mathbf{B}^{*}_{L}$, $\boldsymbol{\beta}^{*}_{1},\ldots,\boldsymbol{\beta}^{*}_{L}$ constitute a saddle-point solution of the max-min problem. Strong duality implies that we can interchange $\min$ and $\max$, i.e., (32) is equivalent to
$\min_{\{\mathbf{R}_{k}\}\in\mathcal{R}}~~\underset{\mathbf{B}_{\ell}\succcurlyeq\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}},\forall\ell}{\mathrm{maximize}}~~\sum_{\ell}2\sqrt{w_{\ell}}\mathbf{e}_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\operatorname*{Tr}(\mathbf{B}_{\ell}\mathbf{J}_{\mathbf{R}}).$ (33)
In particular, for any fixed $\mathbf{R}_{1},\ldots,\mathbf{R}_{K}$, we can optimize over $\mathbf{B}_{\ell}$ by noticing that
$\mathbf{B}_{\ell}\succcurlyeq\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\ \ \Rightarrow\ \ \operatorname*{Tr}(\mathbf{B}_{\ell}\mathbf{J}_{\mathbf{R}})\geq\operatorname*{Tr}(\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{R}}),$ (34)
which holds because $\mathbf{J}_{\mathbf{R}}$ is PSD. Thus, the optimal $\mathbf{B}_{\ell}$ must be $\mathbf{B}_{\ell}=\boldsymbol{\beta}_{\ell}\boldsymbol{\beta}^{\mathsf{T}}_{\ell}$ at the saddle-point. Substituting this relation into (33), and again interchanging maximization and minimization, we arrive at
$\underset{\boldsymbol{\beta}\in\mathbb{R}^{L\times L}}{\mathrm{maximize}}~~\min_{\{\mathbf{R}_{k}\}\in\mathcal{R}}~~\sum_{\ell}\left(2\sqrt{w_{\ell}}\mathbf{e}_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\ \boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{R}}\boldsymbol{\beta}_{\ell}\right),$ (35)
where $\boldsymbol{\beta}\triangleq[\boldsymbol{\beta}_{1},\ldots,\boldsymbol{\beta}_{L}]\in\mathbb{R}^{L\times L}$. This max-min problem is equivalent to the SDR problem (27)), and under Assumption 1, the original ISAC problem (12).
Furthermore, by the following chain of inequalities, the max-min problem (35) can be restated in terms of the beamformers. Let $p^{*}$ denote the optimal value of problem (12).
$p^{*}$ $=\max_{\boldsymbol{\beta}\in\mathbb{R}^{L\times L}}\min_{\{\mathbf{R}_{k}\}\in\mathcal{R}}\sum_{\ell}\left(2\sqrt{w_{\ell}}e_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\ \boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{R}}\boldsymbol{\beta}_{\ell}\right)$ (36a)
$\leq\max_{\boldsymbol{\beta}\in\mathbb{R}^{L\times L}}\min_{\mathbf{V}\in\mathcal{V}}\sum_{\ell}\left(2\sqrt{w_{\ell}}e_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\ \boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{V}}\boldsymbol{\beta}_{\ell}\right)$ (36b)
$\leq\min_{\mathbf{V}\in\mathcal{V}}\max_{\boldsymbol{\beta}\in\mathbb{R}^{L\times L}}\sum_{\ell}\left(2\sqrt{w_{\ell}}e_{\ell}^{\mathsf{T}}\boldsymbol{\beta}_{\ell}-\ \boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{V}}\boldsymbol{\beta}_{\ell}\right)$ (36c)
$=\min_{\mathbf{V}\in\mathcal{V}}\sum_{\ell}w_{\ell}e_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{V}}^{-1}e_{\ell}^{\mathsf{T}}=p^{*},$ (36d)
where (36b) follows from restricting the optimization over $\mathbf{R}_{k}$ to rank-one matrices, (36c) follows from the minimax inequality, and (36d) follows from maximizing the objective in the right-hand side over $\boldsymbol{\beta}$ for fixed $\mathbf{V}$, where the maximum is attained at
$\boldsymbol{\beta}_{\ell}=\sqrt{w}_{\ell}\mathbf{J}_{\mathbf{V}}^{-1}\mathbf{e}_{\ell},\quad\forall\ell.$ (37)
Clearly, each of these inequalities has to be an equality. This establishes the following main result of this section.
Theorem 1
Let $\mathcal{V}$ denote the constraints (12b)-(12c). Suppose the optimization problem (12) is strictly feasible and that Assumption 1 holds. Then, $\mathbf{V}^{*}$ is a global solution of (12) if and only if there exists $\boldsymbol{\beta}^{*}\in\mathbb{R}^{L\times L}$ such that $(\mathbf{V}^{*},\boldsymbol{\beta}^{*})$ is a saddle-point solution of
$\underset{\boldsymbol{\beta}\in\mathbb{R}^{L\times L}}{\mathrm{maximize}}~~\min_{\mathbf{V}\in\mathcal{V}}~~\sum_{\ell=1}^{L}2\sqrt{w_{\ell}}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{e}_{\ell}-\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{V}}\boldsymbol{\beta}_{\ell}$ (38)
where $\boldsymbol{\beta}\triangleq[\boldsymbol{\beta}_{1},\ldots,\boldsymbol{\beta}_{L}]\in\mathbb{R}^{L\times L}$.
Furthermore, problem (38) is equivalent to
$\underset{\boldsymbol{\beta}\in\mathbb{R}^{L\times L}}{\mathrm{maximize}}~\min_{\mathbf{V}\in\mathcal{V}}\sum_{\ell=1}^{L}2\sqrt{w_{\ell}}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{e}_{\ell}-\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{\ell}-\operatorname*{Tr}{\left(\mathbf{Q}_{\boldsymbol{\beta}}\mathbf{V}\mathbf{V}^{H}\right)}$ (39)
where $\mathbf{C}$ is defined in (7) and
$\mathbf{Q}_{\boldsymbol{\beta}}\triangleq\frac{2T}{\sigma^{2}}\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[{(\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{\boldsymbol{(\eta)}})}^{\textsf{H}}\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\right],$ (40)
and
$\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\triangleq\sum_{i}[\boldsymbol{\beta}_{\ell}]_{i}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}$. (41)
Proof:
The development leading to the theorem shows that (12) is equivalent to (38). To further expand the term involving $\mathbf{J}_{\mathbf{V}}$, we note from (6) that
$\sum_{\ell=1}^{L}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{J}_{\mathbf{V}}\boldsymbol{\beta}_{\ell}=\sum_{\ell=1}^{L}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{\ell}+\sum_{\ell=1}^{L}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{T}_{\mathbf{V}}\boldsymbol{\beta}_{\ell},$ (42)
and
$\sum_{\ell=1}^{L}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{T}_{\mathbf{V}}\boldsymbol{\beta}_{\ell}=\sum_{\ell=1}^{L}\sum_{ij}[\boldsymbol{\beta}_{\ell}]_{i}\left[\mathbf{T}_{\mathbf{V}}\right]_{ij}[\boldsymbol{\beta}_{\ell}]_{j}$ (43a)
$=\frac{2T}{\sigma^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\sum_{ij}[\boldsymbol{\beta}_{\ell}]_{i}\mathbb{E}_{\boldsymbol{\eta}}\left[({\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}})^{\textsf{H}}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}[\boldsymbol{\beta}_{\ell}]_{j}\right)$ (43b)
$=\frac{2T}{\sigma^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[\sum_{i}[\boldsymbol{\beta}_{\ell}]_{i}({\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{i}})^{\textsf{H}}\sum_{j}[\boldsymbol{\beta}_{\ell}]_{j}\dot{\mathbf{G}}^{(\boldsymbol{\eta})}_{j}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43c)
$=\frac{2T}{\sigma^{2}}\operatorname*{Tr}\left(\sum_{\ell=1}^{L}\mathbb{E}_{\boldsymbol{\eta}}\left[{(\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{\boldsymbol{(\eta)}})}^{\textsf{H}}\tilde{\mathbf{G}}_{\boldsymbol{\beta_{\ell}}}^{(\boldsymbol{\eta})}\right]\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43d)
$=\operatorname*{Tr}\left(\mathbf{Q}_{\boldsymbol{\beta}}\mathbf{V}\mathbf{V}^{\textsf{H}}\right)$ (43e)
where (43b) follows from (8). ∎
Figure 2: An interpretation of BCRB minimization for the ISAC problem. Different $\boldsymbol{\beta}$ values correspond to different spatial directions. There exists some $\boldsymbol{\beta}^{*}$ for which the minimization of BCRB can be viewed as maximization of power along a spatial direction of interest, subject to the SINR constraints for the communication users and the total power constraint.
Theorem 1 allows us to reformulate the ISAC problem (12) in a more tractable form as the saddle-point solution of a quadratic program as expressed in (39), under Assumption 1. Furthermore, Theorem 1 offers the following important interpretation of BCRB optimization.
Examining the max-min problem (39), by construction the matrix $\mathbf{Q}_{\boldsymbol{\beta}}$ is PSD for any $\boldsymbol{\beta}$. Consequently, the trace term can be interpreted as a sum of beamforming powers measured with respect to the matrix $\mathbf{Q}_{\boldsymbol{\beta}}$. Thus, for a fixed $\boldsymbol{\beta}$, the inner problem in (39) is a power maximization along certain directions of interest. Varying $\boldsymbol{\beta}$ corresponds to varying the directions of interest (as illustrated in Fig. 2). Theorem 1 states that there exists a judicious choice $\boldsymbol{\beta}^{*}$ for which the problem of minimizing the BCRB is equivalent to that of maximizing the power in the direction of $\mathbf{Q}_{\boldsymbol{\beta}^{*}}$.
III-C AoA Estimation Example
Consider again the problem of estimating the AoA for a single target as given in problem (23). For this example, recall that Assumption 1 holds whenever $K\geq 2$. In this case, Theorem 1 can be applied to reformulate problem (23) as a max-min problem.
Since $L=3$, we have $\boldsymbol{\beta}=\left[\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\boldsymbol{\beta}_{3}\right]$, where $\boldsymbol{\beta}_{\ell}\in\mathbb{R}^{3}$. We can further set $\boldsymbol{\beta}_{1}=\boldsymbol{\beta}_{2}=\boldsymbol{0}$ in this case, because the preceding analysis (37) shows that the optimal $\boldsymbol{\beta}_{\ell}$ is given by $\boldsymbol{\beta}_{\ell}^{*}=\sqrt{w_{\ell}}\mathbf{J}_{\mathbf{V}*}^{-1}\mathbf{e}_{\ell}$, and we have $w_{1}=w_{2}=0$, since we are interested in estimating $\theta$ only and not $\alpha$. Then, the set of auxiliary variables reduces to $\boldsymbol{\beta}_{3}=[b_{1},b_{2},b_{3}]^{\mathsf{T}}$ only. By applying Theorem 1, problem (23) can be equivalently expressed as
$\underset{\boldsymbol{\beta}_{3}\in\mathbb{R}^{3}}{\mathrm{maximize}}~\min_{\mathbf{V}\in\mathcal{V}}~2b_{3}-\boldsymbol{\beta}_{3}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{3}-\operatorname*{Tr}{\left(\mathbf{Q}_{\boldsymbol{\beta}_{3}}\mathbf{V}\mathbf{V}^{H}\right)}$ (44)
where
$\mathbf{Q}_{\boldsymbol{\beta}_{3}}=\frac{2T}{\sigma^{2}}\mathbb{E}_{\alpha,\theta}\left[\left(\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right)^{\textsf{H}}\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right$] (45)
and
$\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}}=\left((b_{1}+\imath b_{2})\mathbf{A}(\theta)+b_{3}\alpha\dot{\mathbf{A}}(\theta)\right).$ (46)
We further transform the problem by interchanging max and min, then introducing a new variable $b=\tfrac{(b_{1}+\imath b_{2})}{b_{3}}$, which turns out to allow a simplification of problem (44) as
$\underset{b\in\mathbb{C}}{\mathrm{minimize}}~~\max_{\mathbf{V}\in\mathcal{V}}~~\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\textsf{H}}\right)+\frac{|b|^{2}}{\sigma_{\alpha}^{2}},$ (47)
where $b\in\mathbb{C}$ is the new variable and
$\mathbf{Q}_{b}\triangleq\mathbb{E}_{\alpha,\theta}\left[(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta))^{\textsf{H}}(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta))\right].$ (48)
The equivalence between problem (44) and problem (47) is established in Appendix B by utilizing the fact that $\mathbf{C}\succ 0$, which holds whenever the priors have finite variance.
The above formulations illustrate that the beamforming strategy that minimizes the BCRB also maximizes the power in the direction of some PSD matrix (i.e., the matrix $\mathbf{Q}_{\boldsymbol{\beta}_{3}}\succcurlyeq 0$ in case of problem (44), or the matrix $\mathbf{Q}_{b}\succcurlyeq 0$ in case of problem (47)). This observation is consistent with the interpretation of Theorem 1. Observe that even though we wish to estimate $\theta$ only, the optimal beamforming strategy depends on the distribution of $\alpha$ through the matrices $\mathbf{Q}_{\boldsymbol{\beta}_{3}}$ and $\mathbf{Q}_{b}$.
The expression of $\mathbf{Q}_{b}$ also sheds light on a certain heuristic used in beamforming design for AoA estimation, e.g., [50, 17]. In particular, previous works propose to optimize the signal power, i.e., $\operatorname*{Tr}{(\mathbb{E}_{\theta}[\mathbf{A}(\theta)\mathbf{A}^{\textsf{H}}(\theta)]\mathbf{V}\mathbf{V}^{\textsf{H}})}$, as a heuristic metric to guide the beamforming design. The preceding expression reveals that this heuristic choice, while reasonable, is also not necessarily optimal. Specifically, it shows that optimal direction depends not only on the direction of maximal power as given by the array response $\mathbf{A}(\theta)$, but also on the derivative of $\mathbf{A}(\theta)$ with respect to the parameter of interest $\theta$. Indeed, the optimal direction should depend on some weighted combination of maximizing the signal power in the expected AoA and its sensitivity with respect to the AoA, with weights controlled by the outer minimization in (47).
IV Uplink-Downlink Duality for ISAC
We now proceed to establish an uplink-downlink duality relation for the equivalent formulation (39) of the ISAC problem. This is to say that (39) admits an alternate formulation in terms of a virtual uplink problem in a manner similar to the transformation of the classical downlink communications problem (26) to an uplink problem [10, 12]. However, the duality result for the ISAC problem also differs significantly from the conventional uplink-downlink duality for the communications problem in some important technical aspects as explained below.
IV-A ISAC as Weighted Downlink Power Minimization
We begin by analyzing the inner minimization problem in (39) under the SINR and the total power constraints, for fixed $\boldsymbol{\beta}$, as rewritten explicitly below:
$\underset{\mathbf{v}_{1},\ldots,\mathbf{v}_{K}}{\mathrm{minimize}}$~~~\; $\sum_{\ell=1}^{L}2\sqrt{w_{\ell}}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{e}_{\ell}-\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{\ell}-\operatorname*{Tr}{\left(\mathbf{Q}_{\boldsymbol{\beta}}\mathbf{V}\mathbf{V}^{H}\right)}$ (49a)
$\mathrm{subject\ to}$~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (49b)
$\operatorname*{Tr}\left(\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\leq P.$ (49c)
By dualizing with respect to the total power constraint, the Lagrangian of this optimization problem can be written as
$\mathcal{L}_{\lambda,\boldsymbol{\beta}}\left(\mathbf{V}\right)=\sum_{k=1}^{K}\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}-\lambda P\\ +\sum_{\ell=1}^{L}2\sqrt{w}_{\ell}\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{e}_{\ell}-\boldsymbol{\beta}_{\ell}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{\ell}.$ (50)
The constrained minimization problem (49) can be thought of as a minimization of $\mathcal{L}_{\lambda,\boldsymbol{\beta}}\left(\mathbf{V}\right)$ over $\mathbf{V}$ together with an outer maximization over $\lambda$. Now the maximization over $\lambda$ can be combined with the outer maximization over $\boldsymbol{\beta}$ in (39), so the overall problem (39) can be equivalently written as
$\underset{\boldsymbol{\beta},\lambda\geq 0}{\mathrm{maximize}}~~~\min_{\mathbf{V}}$~~~~~ $\mathcal{L}_{\lambda,\boldsymbol{\beta}}\left(\mathbf{V}\right)$ (51a)
$\mathrm{s.t.\ }$~~~~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k.$ (51b)
This new problem (51) attains the same optimal value as the original ISAC problem (12). Furthermore, at the optimal $(\lambda^{*},\boldsymbol{\beta}^{*})$ of (51), the optimal beamforming strategy of the ISAC problem (12) must be a minimizer of $\mathcal{L}_{\lambda^{*},\boldsymbol{\beta}^{*}}\left(\cdot\right)$. This is because under Assumption 1, the ISAC problem (12) has strong duality (as proved by noting that its SDR is tight, in Lemma 1.)
Now, fix $(\lambda,\boldsymbol{\beta})$ and consider the inner minimization problem in (51). Keeping only the terms involving $\mathbf{v}_{k}$, we have
$\underset{\mathbf{V}}{\mathrm{minimize}}$~~~ $\sum_{k=1}^{K}\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}$ (52a)
$\mathrm{subject\ to\ }$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k.$ (52b)
This problem is now in the same form as the weighted downlink power minimization problem (26) with weight $\mathbf{\tilde{Q}}=\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}$, for which an uplink-downlink duality relationship has been established in [12], except the conventional downlink power minimization problem requires $\mathbf{\tilde{Q}}\succcurlyeq 0$ in (26), but this is not necessarily true for the ISAC problem (52).
This PSD condition on $\mathbf{\tilde{Q}}$ is important because the main result of [12] is that the weight matrix $\mathbf{\tilde{Q}}$ for the downlink power minimization problem becomes the noise covariance matrix in the dual uplink channel. Without the condition $\mathbf{\tilde{Q}}\succcurlyeq 0$, it cannot be a valid noise covariance.
There is no guarantee that for arbitrary $\lambda$ and $\boldsymbol{\beta}$, we would have $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succcurlyeq 0$. Instead, we define the following notion of admissibility to restrict the space of $(\lambda,\boldsymbol{\beta})$ under consideration. It turns out that the optimal $(\lambda^{*},\boldsymbol{\beta}^{*})$ is always admissible, and further, an uplink-downlink duality relationship can be established for all admissible $(\lambda,\boldsymbol{\beta})$’s.
Definition 1
The pair $(\lambda,\boldsymbol{\beta})$ is said to be admissible for a set of channels $(\mathbf{h}_{1},\cdots,\mathbf{h}_{K})$ under the SINR constraints $(\gamma_{1},\cdots,\gamma_{K})$ for the $K$ users with noise variance $\sigma^{2}$, if whenever a set of $\mathbf{V}=[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}]$ satisfies
$\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (53)
we always have
$\sum_{k=1}^{K}\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}\geq 0,\quad\forall k.$ (54)
Admissibility of $(\lambda,\boldsymbol{\beta})$ implies that the matrix $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ behaves somewhat like a PSD matrix, but the condition is much weaker than PSD, not only because $\mathbf{v}_{k}$’s are restricted to the feasibility set satisfying the SINR constraints, but also only the sum power is required to be nonnegative. The individual term $\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}$ can actually be negative.
For inadmissible $(\lambda,\boldsymbol{\beta})$, the minimization problem (52) would be unbounded below, because one can simply scale a set of $\mathbf{v}_{k}$’s for which (54) is negative and drive the objective to negative infinity while satisfying the SINR constraints. Clearly, such $(\lambda,\boldsymbol{\beta})$ cannot be a solution to the outer maximization problem in (51), or equivalently, the original problem (39). For this reason, the search over the optimal $(\lambda,\boldsymbol{\beta})$ in the outer maximization of (51) can be restricted to the admissible set.
More importantly, we can show that whenever $(\lambda,\boldsymbol{\beta})$ is admissible, an uplink-downlink duality result can be established for the inner minimization in (51), or equivalently (52), as explored in detail in the next section. Such uplink-downlink duality would have been a direct consequence of prior work [12] if $\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\succcurlyeq 0$, in which case $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ becomes the noise covariance in the dual uplink channel. The interesting and nontrivial fact is that even if $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is not PSD (so the uplink noise variance may be negative) as long as $(\lambda,\boldsymbol{\beta})$ is admissible, a duality relationship can still be established.
It is worth mentioning that the idea of establishing duality for a downlink problem involving power maximization first appeared in [33] in the context of energy harvesting. The result developed herein is an extension of that in [33]. The admissible set in the setting of [33] is in a one-dimensional space. For ISAC problems, the admissible set is in a much higher $1+L\times L$ dimensional space. This makes the optimization over the admissible set in the ISAC setting much more challenging than the energy harvesting problem.
IV-B Uplink-Downlink Duality
Let $\mathcal{A}$ be the set of admissible $(\lambda,\boldsymbol{\beta})$. The following theorem states an uplink-downlink duality relationship under the condition that $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. The dual uplink channel here has to satisfy an additional technical condition related to a certain matrix being an $M$-matrix.
Theorem 2
Consider the downlink beamforming problem for fixed $(\lambda,\boldsymbol{\beta})$
$\underset{\mathbf{V}}{\mathrm{minimize}}$~~~ $\sum_{k=1}^{K}\mathbf{v}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{v}_{k}$ (55a)
$\mathrm{subject\ to\ }$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (55b)
where the constraints are assumed to be strictly feasible. The optimal value of this downlink optimization problem is bounded below if and only if $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$.
Furthermore, for $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$, an optimal solution $\mathbf{V}^{*}_{\lambda,\boldsymbol{\beta}}\triangleq\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ to (55) can be obtained as
$\mathbf{V}^{*}_{\lambda,\boldsymbol{\beta}}=\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}\mathbf{p}_{\lambda,\boldsymbol{\beta}}^{*},$ (56)
where $\mathbf{U}^{*}_{\lambda,\boldsymbol{\beta}}\in\mathbb{C}^{N_{\text{T}}\times K}$ is a matrix of uplink beamformers obtained by solving the following dual uplink problem over the normalized uplink receive beamformers $\mathbf{U}\triangleq\left[\mathbf{u}_{1},\ldots,\mathbf{u}_{K}\right]$ and the uplink transmit powers $\mathbf{q}\triangleq\left[q_{1},\ldots,q_{K}\right]^{\mathsf{T}}\in\mathbb{R}^{K}$:
$\underset{\mathbf{U},\;\mathbf{q}}{\mathrm{minimize}}$~~ $\sigma^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (57a)
$\mathrm{subject\ to}$~ $\frac{q_{k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}}{\sum_{i\neq k}q_{i}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{k}\right|^{2}+\mathbf{u}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\mathbf{u}_{k}}\geq\gamma_{k},\forall k$ (57b)
$\|\mathbf{u}_{k}\|=1,\quad q_{k}\geq 0,\quad\forall k$ (57c)
$\mathbf{M}_{\mathbf{U}}^{-1}\geq 0,$ (57d)
and $\mathbf{p}^{*}_{\lambda,\boldsymbol{\beta}}\in\mathbb{R}^{K}$ is a vector of downlink powers obtained from
$\mathbf{p}^{*}_{\lambda,\boldsymbol{\beta}}=\sigma^{2}\mathbf{M}_{\mathbf{U}_{\lambda,\boldsymbol{\beta}}^{*}}^{-1}\mathbf{1},$ (58)
where $\mathbf{M}_{\mathbf{U}}\in\mathbb{R}^{K\times K}$ is a matrix defined as
${\left[\mathbf{M}_{\mathbf{U}}\right]_{ij}\triangleq}\left\{\begin{array}[]{cc}\frac{1}{\gamma_{i}}\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{i}\right|^{2},&\mathrm{\ for\quad}i=j,\\ -\left|\mathbf{h}_{i}^{\textsf{H}}\mathbf{u}_{j}\right|^{2},&\mathrm{\ for\quad}i\neq j.\end{array}\right$. (59)
Here, $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0$ in constraint (57d) means that $\mathbf{M}_{\mathbf{U}}$ is invertible and $\mathbf{M}_{\mathbf{U}}^{-1}$ has nonnegative entries.
(a) Downlink beamforming problem for minimizing the transmit power weighted by $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$, while satisfying the downlink SINR constraints.
(b) Uplink beamforming and power minimization problem with noise “covariance” $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ for satisfying the uplink SINR constraints.
Figure 3: Uplink-downlink duality holds for downlink weighted power minimizstion with non-PSD $(\boldsymbol{\lambda}\mathbf{I}-\mathbf{Q}_{\beta})$, if $(\lambda,\boldsymbol{\beta})$ is admissible. The normalized optimal beamformers $\mathbf{v}_{k}$ for the downlink (left) are the same as the optimal beamformers $\mathbf{u}_{k}$ for the uplink (right) with an $M$-matrix constraint.
Proof:
We first prove that the optimal value of the problem (55) is bounded below if and only if $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. Suppose $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$, then by definition of admissibility, for any $\mathbf{V}$ satisfying the SINR constraints, the weighted downlink power is bounded below by zero. Consequently, the optimal value must also be bounded below by zero. Conversely, suppose $(\lambda,\boldsymbol{\beta})\not\in\mathcal{A}$. Then, there exists some $\tilde{\mathbf{V}}$ satisfying the SINR constraints with $\sum_{k}\tilde{\mathbf{v}}_{k}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)\tilde{\mathbf{v}}_{k}<0$. In this case, it can be easily checked that the scaled beamformers $c\tilde{\mathbf{V}}$ with $c>1$ continue to satisfy the SINR constraints. Taking $c\rightarrow\infty$ makes the optimal value unbounded below.
Now consider the downlink problem (55), but over $\mathbf{p}\triangleq\left[p_{1},\ldots,p_{K}\right]^{\mathsf{T}}\in\mathbb{R}^{K}$ and $\mathbf{U}\triangleq\left[\mathbf{u}_{1},\ldots,\mathbf{u}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ with
$\mathbf{u}_{k}=\frac{\mathbf{v}_{k}}{\|\mathbf{v}_{k}\|},\quad p_{k}=\|\mathbf{v}_{k}\|^{2},~~\forall k$ (60)
i.e., $\mathbf{p}$ is a vector of downlink powers and $\mathbf{U}$ is a matrix of normalized beamforming directions. Define the vector ${\boldsymbol{\omega}}_{\mathbf{U}}\in\mathbb{R}^{K}$ as
$\left[{\boldsymbol{\omega}}_{\mathbf{U}}\right]_{k}\triangleq\mathbf{u}_{k}^{\textsf{H}}(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})\mathbf{u}_{k}.$ (61)
Then, we can rewrite problem (55) as a double minimization
$\underset{\mathbf{U}}{\mathrm{minimize}}\quad\underset{\mathbf{p}}{\min}$~~~~ $\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{p}$ (62a)
$\mathrm{s.t.}$~~~~ $\mathbf{M}_{\mathbf{U}}\mathbf{p}\geq\sigma^{2}\mathbf{1},$ (62b)
$\mathbf{p}\geq 0,$ (62c)
$\|\mathbf{u}_{k}\|=1,~\forall k$ (62d)
where (62b) comes from an algebraic manipulation of the SINR constraints (55b).
The matrix $\mathbf{M}_{\mathbf{U}}$ has the specific form of having positive diagonal elements and having off-diagonal elements that are less than or equal to zero. We claim that as long as $\mathbf{U}$, along with some $\mathbf{p}$, is a feasible solution to (62), the matrix $\mathbf{M}_{\mathbf{U}}$ must belong to a class of so-called M-matrices [51]. There are many equivalent definitions of the $M$-matrix; several of these are listed in Appendix C. The reason that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix is because the feasibility of (62b)-(62c) implies the existence of a $\mathbf{p}\geq 0$ for which $\mathbf{M}_{\mathbf{U}}\mathbf{p}>0$. This is one of the equivalent definitions of an $M$-matrix.
The converse to the above statement is also true, i.e., if some $\mathbf{U}$ consisting of normalized beamformers is such that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix, then such $\mathbf{U}$ along with some $\mathbf{p}$ is a feasible solution to (62). The reason such feasible $\mathbf{p}$ can be found is that an $M$-matrix is invertible, and the inverse has nonnegative entries. So, we can obtain a feasible $\mathbf{p}$ by setting (62b) to equality, i.e.,
$\mathbf{p}=\sigma^{2}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{1},$ (63)
which, by the aforementioned property of an $M$-matrix, is always nonnegative, so it automatically satisfies (62c). Thus, $\mathbf{U}$ being feasible is equivalent to $\mathbf{M}_{\mathbf{U}}$ being an $M$-matrix.
We further claim that for any feasible $\mathbf{U}$, the above solution for $\mathbf{p}$ is actually the optimal solution to the inner minimization problem of (62). This would have been a straightforward statement if $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is PSD, in which case the objective of weighted power minimization (62a) would have nonnegative weights, i.e., ${\boldsymbol{\omega}}_{\mathbf{U}}]_{k}\geq 0$, $\forall k$, due to (61). But for this downlink ISAC problem, the weights $\boldsymbol{\omega}_{\mathbf{U}}$ are not necessarily nonnegative. This is because $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ is not necessarily PSD; furthermore, the admissibility condition (54) only ensures that the sum power (or equivalently $\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{p}$) is nonnegative, but allows the individual weights to potentially be negative. This makes sense for sensing problems, because the negative weights indicate the benefit of increasing power toward the target directions.
Nevertheless, we can use the condition $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$ to show that for any feasible $\mathbf{U}$, the inner minimization in (62) would always have the optimal solution achieving equality in (62b), and thus the optimal $\mathbf{p}$ is given by (63). To see this, we make a change of variable and define
$\mathbf{\tilde{p}}\triangleq\mathbf{M}_{\mathbf{U}}\mathbf{p}.$ (64)
Then, the inner minimization in (62) can be rewritten as
$\underset{\mathbf{\tilde{p}}}{\mathrm{minimize}}~$~~~ $\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{\tilde{p}}$ (65a)
$\mathrm{subject\;to}$~~~ $\mathbf{\tilde{p}}\geq\sigma^{2}\mathbf{1},$ (65b)
$\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{\tilde{p}}\geq 0.$ (65c)
We state the following lemma.
Lemma 2
Let $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. Let $\mathbf{U}$ be a matrix of normalized beamformers. Let $\mathbf{M}_{\mathbf{U}}$, $\boldsymbol{\omega}_{\mathbf{U}}$ be defined as in (59) and (61) respectively. Suppose that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix. Then,
$\boldsymbol{\phi}^{\mathsf{T}}_{\mathbf{U}}\triangleq\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{-1}\geq 0.$ (66)
Proof:
By definition, $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$ means that the total downlink power is always nonnegative for any feasible $\mathbf{U}$ and $\mathbf{p}$ satisfying the downlink SINR constraints. Since the objective of the problem (65) is a reformulation of the objective of the downlink beamforming problem (62) or equivalently (55), and $\mathbf{U}$ along with some $\mathbf{p}$ being feasible is equivalent to $\mathbf{M}_{\mathbf{U}}$ being an $M$-matrix, we can state the admissibility condition (54) in terms of the transformed variables $\mathbf{U}$ and $\tilde{\mathbf{p}}$ in the following way: $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$ if and only if the optimal value of (65) is always nonnegative for all $\mathbf{U}$ consisting of normalized beamformers such that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix.
But this is possible only if the coefficients of $\mathbf{\tilde{p}}$ in the objective function (65a), i.e., $\boldsymbol{\phi}^{\mathsf{T}}_{\mathbf{U}}\triangleq\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{-1}$, are all non-negative. This is because of the following. First, problem (65) is clearly feasible. This is because we can set $\mathbf{\tilde{p}}=\mathbf{1}$, which satisfies (65b); further, the fact that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix ensures (65c) is also satisfied. Now, if one of $\boldsymbol{\phi}_{\mathbf{U}}]_{i}$ is negative, then we can arbitrarily increase the corresponding $\mathbf{\tilde{p}}_{i}$ to drive the objective function (65a) to negative without violating the constraints. This is because increasing the value of any $\mathbf{\tilde{p}}_{i}$ clearly does not violate (65b); further, since $\mathbf{M}_{\mathbf{U}}^{-1}$ is the inverse of an $M$-matrix, it has nonnegative entries, so increasing $\mathbf{\tilde{p}}_{i}$ also does not violate the constraint (65c). Since the optimal value of (65) cannot be negative due to the asssumption $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$, we must have $\boldsymbol{\phi}^{\mathsf{T}}_{\mathbf{U}}\geq 0$. ∎
Now, let $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$. For any feasible $\mathbf{U}$ (or equivalently $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix), we have $\boldsymbol{\phi}_{\mathbf{U}}\geq 0$ by Lemma 2. Then, we argue that there must be one optimal solution of (65) that achieves equality in (65b). This is because if the inequality is strict for some $\mathbf{\tilde{p}}_{i}$, one can always decrease $\mathbf{\tilde{p}}_{i}$ without violating (65c) (since $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0$), while improving the objective (65a) or at least keeping it the same (since $\boldsymbol{\phi}^{\mathsf{T}}_{\mathbf{U}}\geq 0$). This shows that $\mathbf{\tilde{p}}=\sigma^{2}\mathbf{1}$ is an optimal solution, so an optimal $\mathbf{p}$ is given by (63).
Plugging this optimal solution $\mathbf{p}$ into the overall downlink beamforming problem (62), we conclude that when $(\lambda,\boldsymbol{\beta})\in\mathcal{A}$, the downlink problem (55) is equivalent to
$\underset{\mathbf{U}}{\mathrm{minimize}}$~~~~ $\sigma^{2}\boldsymbol{\omega}_{\mathbf{U}}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{-1}\mathbf{1}$ (67a)
$\mathrm{subject\;to}$~~~ $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0,$ (67b)
$\|\mathbf{u}_{k}\|=1,~\forall k$ (67c)
where (67b) states that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix. This is needed to ensure that $\mathbf{U}$ is such that a feasible $\mathbf{p}$ exists.
Next, we analyze the uplink problem (57). The idea is also to express the uplink problem as a double minimization
$\underset{\mathbf{U}}{\mathrm{minimize}}\quad\underset{\mathbf{q}}{\min}$~~~~ $\sigma^{2}\mathbf{1}^{\mathsf{T}}\mathbf{q}$ (68a)
$\mathrm{s.t.}$~~~~ $\mathbf{M}_{\mathbf{U}}^{\mathsf{T}}\mathbf{q}\geq\boldsymbol{\omega}_{\mathbf{U}}$ (68b)
$\mathbf{q}\geq 0,\quad\|\mathbf{u}_{k}\|=1,~\forall k$ (68c)
$\mathbf{M}_{\mathbf{U}}^{-1}\geq 0.$ (68d)
For the uplink problem, we choose to explicitly include the requirement that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix as in constraint (68d) to match the downlink formulation. Note that this extra constraint would not have been needed if the uplink noise covariance is PSD, because in that case we have $\boldsymbol{\omega}_{\mathbf{U}}\geq 0$, due to definition (61), and the existence of a feasible power allocation $\mathbf{q}$ implies that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix, due to (68b)-(68c).
As compared to the downlink, the definitions of $\mathbf{M}_{\mathbf{U}}$ and $\boldsymbol{\omega}_{\mathbf{U}}$ here are exactly the same, but the SINR constraints are now expressed in terms of $\mathbf{M}_{\mathbf{U}}^{\mathsf{T}}$. Further, the weights in the power minimization and the noises have their roles reversed.
We now argue that the solution to the inner minimization problem above is again achieved with equality, just as in the downlink. This is because (68b) and (68d) imply that $\mathbf{q}\geq\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$. By Lemma 2, since we assume admissibility, we have that $\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$ must be nonnegative. Now, since the weights in the objective are all positive (i.e., the all-one vector), using the same argument as in the downlink, we conclude that the optimal solution to the inner minimization problem must be
$\mathbf{q}=\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}.$ (69)
Hence, the uplink problem can now be reduced to
$\underset{\mathbf{U}}{\mathrm{minimize}}$~~~~ $\sigma^{2}\mathbf{1}^{\mathsf{T}}\mathbf{M}_{\mathbf{U}}^{\mathsf{-T}}\boldsymbol{\omega}_{\mathbf{U}}$ (70a)
$\mathrm{subject\;to}$~~~ $\mathbf{M}_{\mathbf{U}}^{-1}\geq 0,$ (70b)
$\|\mathbf{u}_{k}\|=1,~\forall k.$ (70c)
Clearly, (70) is the same as (67). By noticing that (58) can be obtained from (63), and (56) is the same as (60), this proves the stated uplink-downlink duality. ∎
Theorem 2 states that whenever the pair $\left(\lambda,\boldsymbol{\beta}\right)$ is admissible, the downlink solution of problem (52) can be recovered by solving the dual uplink problem (57) with $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ as the noise covariance matrix. More specifically, the optimal combiners for the uplink problem (57) are also the optimal beamforming directions for the downlink problem (52) and vice versa. Figure 3 illustrates the duality between the downlink and uplink systems that give rise to problem (52) and problem (57).
The key difference between this new uplink-downlink duality relationship and the classical duality theory is that the uplink noise covariance $(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}})$ here is not necessarily PSD. Thus, there are potentially choices of $\mathbf{u}_{k}$ for which the uplink noise powers are negative. Theorem 2 states that when the additional $M$-matrix constraint is imposed on the uplink beamformers, the uplink-downlink duality continues to hold despite the potentially negative uplink noise powers. This $M$-matrix constraint would not have been needed to establish duality if the uplink noise covariance were PSD.
Appendix D presents a numerical example that shows when the $M$-matrix constraint is omitted from the uplink formulation, there can be choices of $\mathbf{u}_{k}$’s that give rise to feasible solutions for the uplink, but not for the downlink. In this case, the uplink-downlink duality breaks down.
It should be emphasized that the uplink-downlink duality relationship holds only for admissible $(\lambda,\boldsymbol{\beta})$. To illustrate the importance of this admissibility condition, we return to the AoA estimation example.
IV-C AoA Estimation Example
Returning to the example of estimating the AoA while communicating with $K$ users. By incorporating the power constraint into the objective in (47), we obtain the following max-min formulation
$\underset{b\in\mathbb{C},\lambda\geq 0}{\mathrm{maximize}}~\min_{\mathbf{V}}$ $~\operatorname*{Tr}\left(\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{V}\mathbf{V}^{\textsf{H}}\right)-\frac{|b|^{2}}{\sigma_{\alpha}^{2}}-\lambda P$ (71a)
$\mathrm{s.t.}$ $~~\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\sigma^{2}}\geq\gamma_{k},\quad\forall k$ (71b)
For fixed $(\lambda,b)$, the inner problem can be viewed as a downlink communications problem similar to the classical problem (26), provided that the pair $(\lambda,b)$ is admissible.
To check admissibility, we note that the inner problem (71) is bounded if and only if its dual problem (which is a semidefinite program) is feasible. This provides a numerical way of characterizing the admissible set by checking whether the dual problem is feasible for different values of $(\lambda,b)$.
Figure 4: A cut of the admissible set when $\Im{\left\{b\right\}}=1$ for the AoA estimation example with $N_{\text{T}}=N_{\text{R}}=16$, $P=10$, $K=4$ users with SINR constraints set to be $8,10,10$ and 1212dB.
In Figure 4, we show a cut in the admissible set obtained at fixed $\Im\{b\}=1$ across different values of $\lambda$ and $\Re\{b\}$ for a random realization of the channels, where $\alpha\sim\mathcal{CN}(1,1)$ for the sensing channel, and $h_{ij}\sim\mathcal{CN}(0,1)$ for the communication channel. In the figure, the admissible set $\mathcal{A}$ is given by the shaded region. This is to be contrasted with the set of $(\lambda,b)$ values for which $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is PSD, as shown by the dashed curve (which is plotted by setting $\lambda$ to be the largest eigenvalue of $\mathbf{Q}_{b}$).
Theorem 2 states that for any $(\lambda,b)$ outside the admissible set $\mathcal{A}$, the inner minimization in (71) is unbounded. Moreover, for any $(\lambda,b)$ in the admissible region, the inner problem admits an equivalent formulation as an uplink problem with $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ acting as the uplink noise covariance matrix. As seen from the figure, the admissible set $\mathcal{A}$ is strictly larger than the set of $(\lambda,b)$ for which $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is PSD. This underscores the need to develop the new uplink-downlink duality theory for the case where $(\lambda\mathbf{I}-\mathbf{Q}_{b})$ is admissible but not PSD.
If we assume that $b$ is fixed and consider only $\lambda$ for the outer optimization, the admissibility set above reduces to a ray, as in [33] for the energy harvesting scenario. Thus, the above formulation encompasses [33] as a special case [^3] However, the analysis in [33] appears to have overlooked the $M$-matrix constraint for the uplink beamformers and the ensuing complication associated with negative noise power in the duality formulation..
V Numerical Algorithm and Simulations
The uplink-downlink duality is useful in devising low-complexity algorithms for solving the ISAC problem (12). The strategy involves solving the downlink problem (55) for each fixed choice of the pair $(\lambda,\boldsymbol{\beta})$, using its equivalent uplink formulation (57), followed by updating $\left(\lambda,\boldsymbol{\beta}\right)$ in an outer loop. In classical algorithms [10, 9, 11], the uplink problem (57) is solved iteratively. These iterations can be shown to converge to the optimal uplink solution, which then allows the recovery of the optimal downlink solution.
In contrast to the classical case for downlink communications, the duality relation for the ISAC problem only holds for the $(\lambda,\boldsymbol{\beta})$ pairs that are admissible. Furthermore, the uplink problem (57) contains an additional $M$-matrix constraint to ensure the uplink beamformers lead to a feasible solution for the downlink problem.
Despite these complications, it turns out that the classical iterative algorithm of [10, 9, 11] can still be used here, with appropriate modifications, to solve the downlink problem (55) using its uplink counterpart. Furthermore, such a duality algorithm, when combined with a proper outer algorithm for updating $(\lambda,\boldsymbol{\beta})$, can be shown to recover the optimal solution of the ISAC problem (12) under a set of sufficient conditions. The details and the convergence analysis of these algorithms are rather involved and are deferred to a companion paper. Here, we use the AoA estimation example to illustrate the optimized beam patterns obtained from the algorithm.
Consider the recurring example of estimating the AoA for a single target with $N_{\text{T}}=N_{\text{R}}=20$ and $P=10$. For the sensing target, we consider a uniform prior for $\theta$ over the interval $\left[-\Delta,\Delta\right]$, where $\Delta$ is a parameter that determines the range of $\theta$ in the prior. The path loss coefficient $\alpha$ is assumed to follow $\mathcal{CN}(1,1)$ a priori. To avoid the logarithm of the prior to go to minus infinity at the edge of the uniform distribution, a common practice is to approximate the prior by a smooth function or by a Gaussian prior with the same mean and variance for computing the matrix $\mathbf{C}$ in (7); see [48]. Here, we adopt the latter approach for simplicity.
Figure 5: Beam patterns of uplink-downlink duality solution vs. SDR solution. The communication users are at -30 and 50. The sensing target is uniformly distributed at $\Delta,\Delta]$ with $\Delta=5^{\circ}$ or $30^{\circ}$.
For communications, we set $K=2$ users to ensure that Assumption 1 holds. The SINR thresholds are set to be $10$ and $12$ dB. We assume that the communication channels also have a single LoS path with angles at $30^{\circ}$ and $50^{\circ}$ and gain equal to one.
Figure 5 compares the beam pattern of the SDR solution with the beam pattern obtained using the uplink-downlink duality algorithm. The comparison is shown for both narrow and wide priors with $\Delta=5^{\circ}$ and $30^{\circ}$, respectively. In both cases, it is seen that the beam patterns obtained from both algorithms match exactly. Numerically, the optimal values of the objective function also match. Both algorithms achieve a global optimum. The key benefit here is that the duality algorithm is more computationally efficient than SDR, as it does not involve lifting the optimization variables to a higher dimension.
Interestingly, it is further seen from the figure that the optimal beamforming strategy forms beams in the directions of both the communication users and the sensing target, but a wider beam pattern is utilized when $\Delta$ is larger in order to account for the more uncertainty in estimating $\theta$ in that case. We also plot the beam pattern associated with the optimal solution of the classical communication problem (26) without sensing. We see that the resulting beam pattern focuses towards the communication users, but is clearly oblivious of any sensing target, as expected.
VI Conclusion
This paper presents an uplink-downlink duality theory for beamforming design in ISAC systems. We focus on systems for which the use of communication beams alone already attains optimal performance for both communications and sensing, and establish a duality relationship between the downlink beamforming problem and a virtual uplink problem in which the noise covariance matrix is not necessarily PSD. The non-PSD uplink noise covariance arises because in contrast to the conventional downlink beamforming problem for communication alone, the ISAC system aims to maximize the beamformed power in the sensing directions of interest rather than simply to minimize the transmit power. This paper shows that the sensing directions of interest can be characterized precisely via an analysis of the CRB corresponding to the underlying parameter estimation problem. Moreover, the paper establishes an admissibility condition under which the uplink-downlink duality relationship for this new setting holds, and furthermore reveals an additional M-matrix condition that must be included on the virtual uplink beamformers in the duality relationship. These theoretical insights open doors for the development of numerical algorithms for beamformer design that are much more efficient than the conventional SDR approach for downlink beamforming for the ISAC problem.
Appendix A Proof of Lemma 1
Assumption 1 states that the optimal value of problem (12) is equal to the optimal value of the following ISAC optimization that would arise if one instead employs the extended model (2):
$\underset{\mathbf{\tilde{V}}\in\mathbb{C}^{N_{\text{T}}\times(N_{\text{T}}+K)}}{\mathrm{minimize}}~$ $\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}^{-1}_{\mathbf{\tilde{V}}}\right)$ (72a)
$\mathrm{subject\ to}$~~ $\frac{\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{k}\right|^{2}}{\sum_{i\neq k}\left|\mathbf{h}_{k}^{\textsf{H}}\mathbf{v}_{i}\right|^{2}+\mathbf{h}_{k}^{\textsf{H}}\mathbf{V}_{\text{s}}^{\mathsf{H}}\mathbf{V}_{\text{s}}\mathbf{h}_{k}+\sigma^{2}}\geq\gamma_{k},\ \ \forall k$ (72b)
$\|\mathbf{\tilde{V}}\|_{\text{F}}^{2}\leq P$ (72c)
where $\mathbf{\tilde{V}}=\left[\begin{matrix}\mathbf{V}&\mathbf{V}_{\text{s}}\end{matrix}\right]\in\mathbb{C}^{N_{\text{T}}\times(N_{\text{T}}+K)}$ is an extended beamforming matrix with a matrix of communication beamformers $\mathbf{V}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]\in\mathbb{C}^{N_{\text{T}}\times K}$ and a matrix of sensing beamformers $\mathbf{V}_{\text{s}}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{T}}}$. So, to prove the desired result that problem (12) has a tight SDR, we only need to prove that i) problem (72) has a tight SDR; and ii) the SDRs of problems (12) and (72) are equivalent.
The proof of the first claim follows immediately from a standard trick used in several ISAC studies that has been used to show that problems with a similar structure have tight SDR; see, e.g., [4, 6]. We provide the details here to keep this paper self-contained.
Consider the SDR of problem (72), which can be obtained by defining $\mathbf{R}=\mathbf{\tilde{V}}\mathbf{\tilde{V}}^{\mathsf{H}}$, and $\mathbf{R}_{k}=\mathbf{v}_{k}\mathbf{v}_{k}^{\mathsf{H}}$ for all $k$, then dropping the rank-one constraints on $\mathbf{R}_{1},\ldots,\mathbf{R}_{K}$:
$\underset{\mathbf{R},\mathbf{R}_{1},\ldots,\mathbf{R}_{K}}{\mathrm{minimize}}$~~ $\operatorname*{Tr}\left(\mathbf{W}\mathbf{J}_{\mathbf{R}}^{-1}\right)$ (73a)
$\mathrm{subject\ to}$~ $\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}\mathbf{h}_{k}\geq\sigma^{2},\ \ \forall k$ (73b)
$\operatorname*{Tr}\left(\mathbf{R}\right)\leq P,$ (73c)
$\mathbf{R}\succcurlyeq\sum_{k}\mathbf{R}_{k},\quad\mathbf{R}_{k}\succcurlyeq 0~~\forall k.$ (73d)
To show that the above SDR is a tight relaxation of problem (72), we show the existence of a rank-one optimal solution. Let $\mathbf{\hat{R}},\mathbf{\hat{R}}_{1},\ldots,\mathbf{\hat{R}}_{K}$ denote an arbitrary high rank solution of (73). We claim that $\mathbf{R}^{*},\mathbf{R}_{1}^{*},\ldots,\mathbf{R}_{K}^{*}$ as defined below is also a solution:
$\mathbf{R}^{*}=\mathbf{\hat{R}},\quad\mathbf{R}_{k}^{*}=\frac{\mathbf{\hat{R}}_{k}\mathbf{h}_{k}^{\mathsf{H}}\mathbf{h}_{k}\mathbf{\hat{R}}_{k}}{\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{h}_{k}},\quad\forall k.$ (74)
Note that by construction, it is clear that $\mathbf{R}^{*}_{1},\ldots\mathbf{R}^{*}_{K}$ are rank-one PSD matrices.
Moreover, this set of $\mathbf{R}^{*},\mathbf{R}_{1}^{*},\ldots,\mathbf{R}_{K}^{*}$ satisfies the power constraint and attains the same objective value as $\mathbf{\hat{R}},\mathbf{\hat{R}}_{1},\ldots,\mathbf{\hat{R}}_{K}$, because the overall covariance is kept the same, i.e., $\mathbf{R}^{*}=\mathbf{\hat{R}}$.
Next, we note that
$\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{h}_{k}=\mathbf{h}_{k}^{\mathsf{H}}\mathbf{R}^{*}_{k}\mathbf{h}_{k},\quad\forall k$ (75)
which is immediate from (74), so that
$\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\mathsf{H}}\mathbf{R}^{*}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\mathsf{H}}\mathbf{R}^{*}\mathbf{h}_{k}$
$=\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{h}_{k}-\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}\mathbf{h}_{k}\geq\sigma^{2}.$ (76)
In other words, the SINR constraints (73b) are also satisfied.
Finally, for an arbitrary vector $\mathbf{w}\in\mathbb{C}^{N_{\text{T}}}$, the following is true due to Cauchy-Schwarz inequality
$\mathbf{w}^{\mathsf{H}}\left(\mathbf{\hat{R}}_{k}-\mathbf{R}^{*}_{k}\right)\mathbf{w}=\mathbf{w}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{w}-\frac{\left|\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{w}\right|^{2}}{\mathbf{h}_{k}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{h}_{k}}\geq 0.$ (77)
This shows that $\mathbf{w}^{\mathsf{H}}\mathbf{\hat{R}}_{k}\mathbf{w}\geq\mathbf{w}^{\mathsf{H}}\mathbf{R}^{*}_{k}\mathbf{w}$, $\forall\mathbf{w}$, or equivalently $\mathbf{\hat{R}}_{k}\succcurlyeq\mathbf{R}^{*}_{k}$, which implies that
$\mathbf{R}^{*}=\mathbf{\hat{R}}\succcurlyeq\sum_{k}\mathbf{\hat{R}}_{k}\succcurlyeq\sum_{k}\mathbf{R}^{*}_{k}.$ (78)
In other words, the constraint (73d) is also satisfied. This shows that the SDR (73) has a rank-one solution and that the relaxation is tight.
Next, we aim to prove that the SDR of problem (12) (as given in (27)), and the SDR problem (73) are equivalent. To do so, it suffices to show that any feasible solution of one problem can be mapped into a feasible solution for the other problem. Let $\mathbf{\tilde{R}}_{1},\ldots,\mathbf{\tilde{R}}_{K}$ be an arbitrary feasible solution of the SDR (27). Then, it is easily verified that the solution $\mathbf{R},\mathbf{R}_{1},\ldots,\mathbf{R}_{K}$ defined below satisfies all constraints in (73)
$\mathbf{R}=\sum\mathbf{\tilde{R}}_{k},\quad\mathbf{R}_{k}=\mathbf{\tilde{R}}_{k},\quad\forall k.$ (79)
Conversely, suppose that $\mathbf{R},\mathbf{R}_{1},\ldots,\mathbf{R}_{K}$ is a feasible solution of problem (73). Then, we claim that
$\mathbf{\tilde{R}}_{k}=\mathbf{R}_{k}+\frac{1}{K}\left(\mathbf{R}-\sum_{i}\mathbf{R}_{i}\right),\quad\forall k$ (80)
is a feasible solution of (27). To see this, we note that
$\mathbf{\tilde{R}}_{k}\succcurlyeq\mathbf{R}_{k}\succcurlyeq 0,\quad\forall k$ (81)
since the second term in (80) is PSD as given by (73d). Next, we note that
$\sum_{k}\mathbf{\tilde{R}}_{k}=\sum_{k}\left(\mathbf{R}_{k}+\frac{1}{K}\left(\mathbf{R}-\sum_{i}\mathbf{R}_{i}\right)\right)$ (82a)
$=\left(\sum_{k}\mathbf{R}_{k}\right)+\mathbf{R}-\sum_{i}\mathbf{R}_{i}=\mathbf{R}$ (82b)
i.e., the matrices $\mathbf{\tilde{R}}_{k}$ in (80) sum up to $\mathbf{R}$. Therefore, the power constraint in (27) is satified for $\mathbf{\tilde{R}}_{1},\ldots,\mathbf{\tilde{R}}_{K}$, i.e.,
$\operatorname*{Tr}\left(\sum_{k}\mathbf{\tilde{R}}_{k}\right)=\operatorname*{Tr}(\mathbf{R})\leq P.$ (83)
Finally, the SINR constraints in (27) (i.e., constraints (27b)) are also satisfied due to the following
$\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\textsf{H}}\mathbf{\tilde{R}}_{k}\mathbf{h}_{k}$ $\geq\left(1+\frac{1}{\gamma_{k}}\right)\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}_{k}\mathbf{h}_{k}$ (84a)
$\geq\mathbf{h}_{k}^{\textsf{H}}\mathbf{R}\mathbf{h}_{k}+\sigma^{2}$ (84b)
$=\sum_{k}\mathbf{h}_{k}^{\textsf{H}}\mathbf{\tilde{R}}_{k}\mathbf{h}_{k}+\sigma^{2}$ (84c)
where the first line follows from (81), the second line follows from the SINR constraint (73b), and the third line follows from (82). This shows that SDRs (27) and (73) are equivalent.
Combining the above with Assumption 1 and the fact that the relaxation (73) is tight, we conclude that all four problems (i.e., problem (12), problem (72), and their SDRs) are equivalent.
Appendix B Equivalence of Problem (44) and Problem (47)
First, recall that we assume that $\alpha$ and $\theta$ are independent with Gaussian priors so that
$\mathbf{C}=\frac{2T}{\sigma^{2}}\operatorname*{diag}\left(\frac{1}{\sigma^{2}_{\alpha}},\frac{1}{\sigma^{2}_{\alpha}},\frac{1}{\sigma^{2}_{\theta}}\right)\succ 0.$ (85)
It follows that $\mathbf{J}_{\mathbf{V}}\succ 0$ in (18), because $\mathbf{J}_{\mathbf{V}}=\mathbf{C}+\mathbf{T}_{\mathbf{V}}$ and $\mathbf{T}_{\mathbf{V}}\succcurlyeq 0$.
Under Assumption 1, we can interchange the $\max$ and the $\min$ in problem (44) to obtain the equivalent problem
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{minimize}}~\max_{b_{1},b_{2},b_{3}}$~ $2b_{3}-\boldsymbol{\beta}_{3}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{3}-\operatorname*{Tr}{\left(\mathbf{Q}_{\boldsymbol{\beta}_{3}}\mathbf{V}\mathbf{V}^{H}\right)}.$ (86)
We now claim that the optimal $b_{3}$ in the inner problem above always satisfy $b_{3}>0$. This is because we can use (42), (43) and $b_{3}=\mathbf{e}_{3}^{\mathsf{T}}\boldsymbol{\beta}_{3}$ to rewrite the inner problem as
$\underset{\boldsymbol{\beta}_{3}\in\mathbb{R}^{3}}{\mathrm{maximize}}~~2\mathbf{e}_{3}^{\mathsf{T}}\boldsymbol{\beta}_{3}-\boldsymbol{\beta}_{3}^{\mathsf{T}}\mathbf{J}_{\mathbf{V}}\boldsymbol{\beta}_{3},$ (87)
which has a solution $\boldsymbol{\beta}_{3}^{*}=\mathbf{J}^{-1}_{\mathbf{V}}\mathbf{e}_{3}$. Noting that $\mathbf{J}_{\mathbf{V}}\succ 0$, this implies
$b_{3}^{*}=\mathbf{e}_{3}^{\mathsf{T}}\boldsymbol{\beta}_{3}^{*}=\mathbf{e}_{3}^{\mathsf{T}}\mathbf{J}^{-1}_{\mathbf{V}}\mathbf{e}_{3}>0.$ (88)
Thus, we can restrict to $b_{3}>0$ without loss of optimality.
Now, consider the following change of variables:
$a\triangleq b_{3}\in\mathbb{R}_{++},\quad b\triangleq\tfrac{1}{b_{3}}(b_{1}+\imath b_{2})\in\mathbb{C},$ (89)
which is related to the perspective function (see [45]).
Since $b_{3}\neq 0$, the mapping is well defined. Furthermore, given $(a,b)$ with $a>0$, we can uniquely recover $b_{1}=\Re\{ab\}$, $b_{2}=\Im\{ab\}$, and $b_{3}=a$. In fact, the mapping from $(b_{1},b_{2},b_{3})\in\mathbb{R}^{2}\times\mathbb{R}_{++}$ to $(a,b)\in\mathbb{R}_{++}\times\mathbb{C}$ is bijective. This means that instead of optimizing over $(b_{1},b_{2},b_{3})$, we can reformulate the problem as an optimization over $(a,b)$.
Now, it can be verified that
$\boldsymbol{\beta}_{3}^{\mathsf{T}}\mathbf{C}\boldsymbol{\beta}_{3}$ $=\frac{2T}{\sigma^{2}\sigma^{2}_{\alpha}}\left(b_{1}^{2}+b_{2}^{2}\right)+\frac{2T}{\sigma^{2}\sigma^{2}_{\theta}}b^{2}_{3}$ (90a)
$=\frac{2T}{\sigma^{2}}a^{2}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}\right),$ (90b)
and $\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}}=(b_{1}+\imath b_{2})\mathbf{A}(\theta)+b_{3}\alpha\dot{\mathbf{A}}(\theta)=a(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta))$, so that
$\mathbf{Q}_{\boldsymbol{\beta}_{3}}$ $=\frac{2T}{\sigma^{2}}\mathbb{E}_{\alpha,\theta}\left[\left(\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right)^{\textsf{H}}\tilde{\mathbf{G}}^{(\alpha,\theta)}_{\boldsymbol{\beta}_{3}}\right$] (91a)
$=\frac{2T}{\sigma^{2}}a^{2}\mathbb{E}_{\alpha,\theta}\left[\left(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta)\right)^{\textsf{H}}\left(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta)\right)\right]$ (91b)
$=\frac{2T}{\sigma^{2}}a^{2}\mathbf{Q}_{b}.$ (91c)
where $\mathbf{Q}_{b}$ is defined in (48). Then, problem (44) can be stated as
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{minimize}}\max_{a>0,b}2a-\frac{2T}{\sigma^{2}}a^{2}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right)$ (92)
For fixed $\mathbf{V}$ and $b$, the maximization over $a$ occurs at
$a^{*}=\frac{1}{\frac{2T}{\sigma^{2}}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right)}>0.$ (93)
Substituting back into (92) and taking the reciprocal yield
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{maximize}}~\min_{b\in\mathbb{C}}$ $~~\frac{2T}{\sigma^{2}}\left(\frac{|b|^{2}}{\sigma_{\alpha}^{2}}+\frac{1}{\sigma^{2}_{\theta}}+\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right)\right),$ (94)
As final step, we eliminate the constants $\tfrac{2T}{\sigma^{2}}$ and $\frac{1}{\sigma^{2}_{\theta}}$, and again interchange max and min to arrive at the desired form (47).
It turns out that the alternative formulation (47) can also be derived by analyzing the dual of the following equivalent formulation of problem (23):
$\underset{\mathbf{V}\in\mathcal{V}}{\mathrm{maximize}}$~~ $d^{\prime}$ (95a)
$\mathrm{subject\;to}$~~ $\left[\begin{matrix}h_{\theta}(\mathbf{V})-d^{\prime}&h_{\alpha\theta}(\mathbf{V})\\ \bar{h}_{\alpha\theta}(\mathbf{V})&h_{\alpha}(\mathbf{V})+\tfrac{1}{\sigma_{\alpha}^{2}}\end{matrix}\right]\succcurlyeq 0.$ (95b)
This can be accomplished by dualizing with respect to constraint (95b) and by following the same steps as in the proof of Theorem 1.
Let $\mathbf{\tilde{B}}$ denote the $2\times 2$ dual matrix associated with (95b):
$\mathbf{\tilde{B}}=\left[\begin{matrix}b^{\prime}&\bar{b}\\ b&b^{\prime\prime}\end{matrix}\right]\succcurlyeq 0$ (96)
where $^{\prime}\in\mathbb{R}, bb\in\mathbb{C}$, and $b^{\prime\prime}\in\mathbb{R}$. The Lagrangian dual of the problem (95) is
$\underset{\mathbf{\tilde{B}}\succcurlyeq 0}{\mathrm{minimize}}\max_{\mathbf{V}\in\mathcal{V},d^{\prime}}d^{\prime}(1-b^{\prime})+b^{\prime}h_{\theta}(\mathbf{V})+\bar{b}\bar{h}_{\alpha\theta}(\mathbf{V})$
$~~~~+bh_{\alpha\theta}(\mathbf{V})+b^{\prime\prime}\left(h_{\alpha}(\mathbf{V})+\tfrac{1}{\sigma^{2}_{\alpha}}\right)$
which is equivalent to problem (23) due to strong duality (i.e., under Assumption 1). In the above problem, we must have $b^{\prime}=1$ since otherwise the inner optimization is unbounded. Substituting $b^{\prime}=1$, we get
$\underset{b^{\prime\prime}\geq|b|^{2}}{\mathrm{minimize}}~\max_{\mathbf{V}\in\mathcal{V}}h_{\theta}(\mathbf{V})+\bar{b}\bar{h}_{\alpha\theta}(\mathbf{V})+bh_{\alpha\theta}(\mathbf{V})$
$+b^{\prime\prime}\left(h_{\alpha}(\mathbf{V})+\tfrac{1}{\sigma^{2}_{\alpha}}\right)$ (97)
where the condition $b^{\prime\prime}\geq|b|^{2}$ follows from $\tilde{\mathbf{B}}\succcurlyeq 0$ and $b^{\prime}=1$, so the Schur complement must satisfy $b^{\prime\prime}-|b|^{2}\geq 0$.
By interchanging min and max, it is easy to see that the optimal $b^{\prime\prime}$ is attained when $b^{\prime\prime}=|b|^{2}$, since $(h_{\alpha}(\mathbf{V})+\tfrac{1}{\sigma^{2}_{\alpha}})$ is positive for every $\mathbf{V}$. Substituting this into (B), we arrive at
$\underset{b\in\mathbb{C}}{\mathrm{minimize}}~\max_{\mathbf{V}\in\mathcal{V}}~~h_{\theta}(\mathbf{V})+\bar{b}\bar{h}_{\alpha\theta}(\mathbf{V})+bh_{\alpha\theta}(\mathbf{V})$
$+|b|^{2}h_{\alpha}(\mathbf{V})+\tfrac{|b|^{2}}{\sigma^{2}_{\alpha}}$ (98)
Finally, it can be verified that:
$h_{\theta}(\mathbf{V})+\bar{b}\bar{h}_{\alpha\theta}(\mathbf{V})+bh_{\alpha\theta}(\mathbf{V})\\ +|b|^{2}h_{\alpha}(\mathbf{V})=\operatorname*{Tr}\left(\mathbf{Q}_{b}\mathbf{V}\mathbf{V}^{\mathsf{H}}\right),$
so that problem (B) is precisely problem (47).
Appendix C Brief Review of $M$-Matrices
A matrix $\mathbf{B}$ is termed an $M$-matrix if two conditions hold:
  • The diagonal elements are nonnegative but the off-diagonal elements are nonpositive, i.e.,
    $\mathbf{B}]_{ii}\geq 0,~\forall i.\quad\text{and}\quad[\mathbf{B}]_{ij}\leq 0,~\forall i\neq j,$ (99)
  • The inverse exists and has nonnegative elements, i.e., $\mathbf{B}^{-1}]_{ij}\geq 0,\forall i,j$.
$M$-matrices play an important role in stability analysis in Control Theory. For this reason, they have been extensively studied. There exist several equivalent characterizations of the second condition, as long as the first condition above holds. Of interest to this paper are the following equivalent conditions:
  1. 1.
    $\mathbf{B}^{-1}$ exists and is nonnegative.
  2. 2.
    There exists a vector $\mathbf{p}\geq 0$ satisfying $\mathbf{B}\mathbf{p}>0$.
  3. 3.
    $\mathbf{B}=\mathbf{D}-\mathbf{F}$ with $\mathbf{D}^{-1}\>{\geq}~0$, $\mathbf{F}\geq 0$ and $\rho_{\text{max}}(\mathbf{D}^{-1}\mathbf{F})<1$.
A proof of equivalence of these conditions, as well as an extensive list of equivalent characterizations can be found in [51].
Appendix D Necessity of $M$-Matrix Constraint in the Dual Uplink Formulation: An Example
We use a numerical example to demonstrate that when $\left(\lambda\mathbf{I}-\mathbf{Q}_{\boldsymbol{\beta}}\right)$ is not PSD, the constraint that $\mathbf{M}_{\mathbf{U}}$ is an $M$-matrix must be included in the uplink formulation (57) for the uplink-downlink duality relationship to hold. This is so even for admissible $(\lambda,\boldsymbol{\beta})$.
Let $N_{\text{T}}=N_{\text{R}}=2$ and $K=2$ with the following channel realizations
$\mathbf{h}_{1}=\left[\begin{matrix}1\\ 0\end{matrix}\right],\quad\mathbf{h}_{2}=\left[\begin{matrix}0\\ 1\end{matrix}\right],$ (100)
and the following SINR thresholds and noise power:
$\gamma_{1}=4,\quad\gamma_{2}=2,\quad\sigma^{2}=1.$ (101)
For sensing, we consider the AoA estimation problem with independent $\alpha$ and $\theta$. The priors are $\alpha\sim\mathcal{CN}(0,1)$ and $\theta\sim\mathcal{N}(0,1)$. Set $b=1$ as in (48). We can calculate
$\mathbf{Q}_{b}$ $=\mathbb{E}_{\alpha,\theta}\left[(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta))^{\textsf{H}}(b\mathbf{A}(\theta)+\alpha\dot{\mathbf{A}}(\theta))\right]$ (102a)
$\approx\left[\begin{matrix}1.2566&-0.0458\\ -0.0458&1.2566\end{matrix}\right].$ (102b)
This matrix has equal diagonal elements and has eigenvectors
$\mathbf{w}_{1}=\frac{1}{\sqrt{2}}\left[\begin{matrix}1\\ -1\end{matrix}\right],\quad\mathbf{w}_{2}=\frac{1}{\sqrt{2}}\left[\begin{matrix}1\\ 1\end{matrix}\right],$ (103)
and eigenvalues
$\rho_{1}=1.3024,\quad\rho_{2}=1.2108.$ (104)
The set of admissible $(\lambda,b)$’s are those for which the downlink problem is bounded below:
$\underset{\mathbf{v}_{1},\mathbf{v}_{2}}{\mathrm{minimize}}$~~ $\mathbf{v}_{1}^{\mathsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{v}_{1}+\mathbf{v}_{2}^{\mathsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{v}_{2}$ (105a)
$\mathrm{subject\ to}$~ $\frac{\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{v}_{1}\right|^{2}}{\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{v}_{2}\right|^{2}+1}\geq 4,\quad\frac{\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{v}_{2}\right|^{2}}{\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{v}_{2}\right|^{2}+1}\geq 2.$ (105b)
This problem has strong duality. One way to check whether (105) is bounded below is to check whether its dual problem, as given below, is feasible:
$\underset{z_{1}\geq 0,z_{2}\geq 0}{\mathrm{minimize}}$~~ $z_{1}+z_{2}$ (106a)
$\mathrm{subject\ to}$~~ $\lambda\mathbf{I}-\mathbf{Q}_{b}+z_{2}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}\succcurlyeq\tfrac{z_{1}}{4}\mathbf{h}_{1}\mathbf{h}_{1}^{\mathsf{H}},$ (106b)
$\lambda\mathbf{I}-\mathbf{Q}_{b}+z_{1}\mathbf{h}_{1}\mathbf{h}_{1}^{\mathsf{H}}\succcurlyeq\tfrac{z_{2}}{2}\mathbf{h}_{2}\mathbf{h}_{2}^{\mathsf{H}}.$ (106c)
Numerically, one can determine that the minimum value of $\lambda$ for which $(\lambda,b)$ is admissible is $\lambda_{\text{min}}\approx 1.297$.
Now, choose
$\lambda=1.3>\lambda_{\text{min}},$ (107)
so that the pair $(\lambda,b)=(1.3,1)$ is admissible. Observe that,
$\lambda\mathbf{I}-\mathbf{Q}_{b}=\left[\begin{matrix}0.0434&0.0458\\ 0.0458&0.0434\end{matrix}\right]$ (108)
is not PSD since $\lambda<\rho_{1}$.
For this $(\lambda\mathbf{I}-\mathbf{Q}_{b})$, we claim that the $M$-matrix constraint cannot be dropped from the uplink formulation, otherwise the duality relationship in Theorem 2 would not hold. To see this, we first solve the downlink problem (105) numerically by SDR. The optimal value of this problem is approximately
$f_{\text{DL}}^{*}\approx 0.1435.$ (109)
Now, consider the uplink problem as stated below without the $M$-matrix constraint:
$\underset{q_{1},q_{2},\mathbf{u}_{1},\mathbf{u}_{2}}{\mathrm{minimize}}$~~ $q_{1}+q_{2}$ (110a)
$\mathrm{subject\ to}$~ $\frac{q_{1}\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{u}_{1}\right|^{2}}{q_{2}\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{u}_{1}\right|^{2}+\mathbf{u}_{1}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{1}}\geq 4$ (110b)
$\frac{q_{2}\left|\mathbf{h}_{2}^{\textsf{H}}\mathbf{u}_{2}\right|^{2}}{q_{1}\left|\mathbf{h}_{1}^{\textsf{H}}\mathbf{u}_{2}\right|^{2}+\mathbf{u}_{2}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{2}}\geq 2$ (110c)
$q_{1}\geq 0,\quad q_{2}\geq 0.$ (110d)
We proceed to show that the optimal value of the above problem is strictly less than $f_{\text{DL}}^{*}$
Set the receive beamformers in the uplink problem to be the eigenvector corresponding to $\rho_{1}$, i.e., $\mathbf{u}_{1}=\mathbf{u}_{2}=\mathbf{w}_{1}$. This choice of beamformers may not be optimal for the uplink problem. However, we show next that such a choice leads to a feasible power allocation that already achieves a strictly lower objective value for the uplink problem than $f_{\text{DL}}^{*}$.
For the above choice, the “virtual” uplink noise powers are negative
$\mathbf{u}_{1}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{1}=\mathbf{u}_{2}^{\textsf{H}}\left(\lambda\mathbf{I}-\mathbf{Q}_{b}\right)\mathbf{u}_{2}\approx-0.0024,$ (111)
and all the direct and interfering channels are equal, i.e.,
$|\mathbf{h}_{i}^{\mathsf{H}}\mathbf{u}_{j}|^{2}=\tfrac{1}{2},\quad\forall i,j=1,2.$ (112)
With this choice of $({\mathbf{u}}_{1},{\mathbf{u}}_{2})$, the uplink constraints (110b)-(110d) simplify to:
$\frac{\tfrac{1}{2}q_{1}}{\tfrac{1}{2}q_{2}-0.0024}\geq 4,~~~\frac{\tfrac{1}{2}q_{2}}{\tfrac{1}{2}q_{1}-0.0024}\geq 2,~~~q_{1}\geq 0,~~~q_{2}\geq 0.$ (113)
The following choice of $(q_{1},q_{2})$ satisfies the above set of inequalities:
$q_{1}=q_{2}\approx\frac{8}{3}\cdot 0.0024.$ (114)
Furthermore, they give a sum power less than $f_{\text{DL}}^{*}$, i.e., we have
$f_{\text{UL}}^{*}\leq q_{1}+q_{2}<f_{\text{DL}}^{*}.$ (115)
In other words, the downlink problem (105) and the uplink problem (110) have different optimal values. Therefore, the uplink-downlink duality relationship does not hold.
To further confirm the importance of the $M$-matrix constraint, we demonstrate that the choice $\mathbf{u}_{1}=\mathbf{u}_{2}=\mathbf{w}_{1}$ does not give rise to a feasible solution for the downlink problem (105), despite being a feasible choice for the uplink problem (110).
Set $\mathbf{v}_{1}=\sqrt{p_{1}}\mathbf{w}_{1}$ and $\mathbf{v}_{2}=\sqrt{p_{2}}\mathbf{w}_{1}$ as the downlink beamformers in the SINR constraint (105b). This yields
$\frac{\tfrac{1}{2}p_{1}}{\tfrac{1}{2}p_{2}+1}\geq 4,\quad\frac{\tfrac{1}{2}p_{2}}{\tfrac{1}{2}p_{1}+1}\geq 2,$ (116)
which have no solution for $p_{1}\geq 0$ and $p_{2}\geq 0$. Indeed, this is because the above constraints can be written as:
$\left[\begin{matrix}\tfrac{1}{8}&-\tfrac{1}{2}\\ -\tfrac{1}{2}&\tfrac{1}{4}\end{matrix}\right]\left[\begin{matrix}p_{1}\\ p_{2}\end{matrix}\right]\geq\left[\begin{matrix}1\\ 1\end{matrix}\right],$ (117)
and the matrix on the right-hand side has an inverse with all-negative entries, so it is not an $M$-matrix. Multiplying both sides by the matrix inverse and flipping the inequality, we have
$\left[\begin{matrix}p_{1}\\ p_{2}\end{matrix}\right]\leq\left[\begin{matrix}-\tfrac{24}{7}\\ -\tfrac{20}{7}\end{matrix}\right].$ (118)
Since the power cannot be negative, this choice of beamformers is not feasible for the downlink problem.
As can be seen in this example, it is the negative noise power in the uplink channel that allows the set of feasible beamformers in the uplink to be larger than the corresponding set of feasible beamformers in the downlink. The purpose of the $M$-matrix constraint is to restrict the feasible set of uplink beamformers to be exactly those beamformers that are also feasible in the downlink. This is a crucial constraint that cannot be omitted.